### v2

# Spaces of subgroups of toral groups

Abstract.
We study the space of conjugacy classes of subgroups of a compact Lie group $G$ whose identity component is a torus, and consider how various invariants of subgroups behave as sheaves over this space. This feeds in to the author’s programme to give algebraic models of rational $G$-equivariant cohomology theories.
The author is grateful for comments, discussions and related collaborations with S.Balchin, D.Barnes, T.Barthel, Z.Huang, M.Kedziorek, L.Pol, J.Williamson. The work is partially supported by EPSRC Grant EP/W036320/1. The author would also like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programmes Topology, representation theory and higher structures and Equivariant homotopy theory in context, where later parts of work on this paper was undertaken. This work was supported by EPSRC grant EP/Z000580/1. The author declares there are no competing interests.
1. Introduction
1.A. Contents
The paper is part of a general programme to understand equivariant cohomology theories. We explain this briefly for the interested reader in Subsection 1.B below, but the mathematical content of the present paper is purely group theoretical.
This paper is about the space $\mathfrak{X}_{G}=\mathrm{Sub}(G)/G$ of conjugacy classes of (closed) subgroups of a compact Lie group $G$. More specifically it considers the special case that $G$ is a finite extension of a torus:
$1\longrightarrow\mathbb{T}\longrightarrow G\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}W\longrightarrow 1.$
It gives a method for identifying the set $\mathfrak{X}_{G}$ as well as certain additional structure, including the Hausdorff metric topology and the normalizers of subgroups.
The group $G$ is determined by an action of $W$ on $\mathbb{T}$ and a cohomology class determining the extension. Since $\mathrm{Aut}(\mathbb{T})=\mathrm{Aut}(H_{1}(\mathbb{T};\mathbb{Z}))$, the action is specified by the $\mathbb{Z}W$-module $\Lambda_{0}=H_{1}(\mathbb{T};\mathbb{Z})$, which we call the toral latice . We then write $L\mathbb{T}=\Lambda_{0}\otimes\mathbb{R}$ for the associated real representation; by virtue of the exponential map $\mathbb{T}=L\mathbb{T}/\Lambda_{0}$. The extension class $\epsilon(G)$ lies in $H^{2}(W;\mathbb{T})\cong H^{3}(W;\Lambda_{0})$.
In effect, the identification of $\mathfrak{X}_{G}$ is just an extended exercise exploiting Pontrjagin duality, which states that subgroups of the compact Lie group $\mathbb{T}$ correspond to subgroups of the discrete abelian group $\mathbb{T}^{*}=\mathrm{Hom}(\mathbb{T},T)$, where $T$ is the circle group. The point is to do it in such a way that we keep track of the dependence of invariants on subgroups.
We partition the subgroups $K$ according to the conjugacy class of the image of $K$ in $W$, and for the most part, we consider only ‘full’ subgroups (those mapping onto $W$). To any full subgroup $K$ we associate the subgroup $S=K\cap\mathbb{T}$, which is invariant under the conjugation action of $W$. This gives a map
$c:\mathfrak{X}^{\mathrm{full}}_{G}\longrightarrow\mbox{$W$-$\mathrm{Sub}(\mathbb{T})$}\cong\mbox{$W$-$\mathrm{Sub}(\mathbb{T}^{*})$},$
and we show (Corollary 3.5) that $c$ is an isomorphism up to finite indeterminacy: it is almost surjective, and has finite fibres. There are finitely many different fibres, and these are cohomologically determined. We can also understand how other invariants of subgroups (such as the Weyl groups $W_{G}(K)$) vary with $S$.
When working out $\mathfrak{X}_{G}$ it is helpful to have already worked out $\mathfrak{X}_{H}$ for subgroups $H$ of $G$, but some modifications are necessary (i) several $H$-conjugacy classes may fuse to form a $G$-conjugacy class, so the space $\mathfrak{X}_{G}|H$ is a quotient of $\mathfrak{X}_{H}$ and (ii) the normalizer of a subgroup $K$ in $H$ is a subgroup of the normalizer in $G$ and hence the $H$-Weyl group $W_{H}(K)$ is a subgroup of the $G$-Weyl group $W_{G}(K)$.
We show that the method is effective by giving explicit descriptions for a number of rank 2 groups $G$. The reader wondering if this apparatus is necessary might wish to try a direct approach on a simple example. The case of the normalizer of the maximal torus in $U(2)$ (which is nothing more than the semidirect product $T^{2}\rtimes C_{2}$ where $C_{2}$ exchanges the factors) is instructive. The direct approach is certainly possible, but the apparatus helps organize and understand the answer.
1.B. Motivation
We briefly explain the larger project of which this is a part.
For a compact Lie group $G$, it has been conjectured [6] that the category of rational $G$-spectra is Quillen equivalent to differential graded objects of an abelian category $\mathcal{A}(G)$. The abelian model $\mathcal{A}(G)$ is a category of sheaves over the space $\mathfrak{X}_{G}=\mathrm{Sub}(G)/G$ of conjugacy classes of subgroups of $G$. The basic data associated to a subgroup $K$ is a module over the polynomial ring $H^{*}(BW_{G}^{e}(K))$ with an action of the finite group $W_{G}^{d}(K)$ where $W_{G}(K)=N_{G}(K)/K$ is the Weyl group with identity component $W_{G}^{e}(K)$ and component group $W_{G}^{d}(K)$. As the subgroup varies, $H^{*}(BW_{G}^{e}(K))$ gives a ‘sheaf’ of rings, and $W_{G}^{d}(K)$ gives a ‘component structure’. The basic input data is therefore the space $\mathrm{Sub}(G)/G$ together with the Weyl group $W_{G}(K)$ for each subgroup $K$.
[…]
1.C. Organization
In Part 1, we give general methods for classifying the subgroups of toral groups $G$, and identifying the properties that feed in to an algebraic model. The work begins in Section 2, by preparing the ground for the classification of subgroups: partitioning them according to subgroups of the component group and setting out the use of Pontrjagin duality. Section 3 gives the classification of subgroups up to conjugacy in cohomological terms. Section 4 explains how to use linear algebra to calculate normalizers and Weyl groups through lattices. In Section 5 we describe the topology on the space of subgroups in terms of dual lattices.
In Part 2, we apply the methods to all toral groups occuring as subgroups of a rank 2 connected compact Lie group. We begin in Section 6 by summarising the method. The cases treated are $A_{1}\times A_{1}$ (with component group $W=D_{4}$ or a subgroup), $A_{2}$ (with group component group $W=D_{6}$ or a subgroup), $B_{2}$ (with component group $W=D_{8}$ or a subgroup) or $G_{2}$ (with component group $W=D_{12}$ or a subgroup). This makes for a large number of cases, and they are listed in a table in Section 6. The discussion begins in Section 7 with a summary of the group cohomology calculations required for all examples. This is a routine calculation, but it is convenient to have it layed out for reference. After this in Sections 8 to 17 we treat the examples in turn (the table in Subsection 6.B points to the exact section where each is discussed).
1.D. Associated work in preparation
This paper describes the group theoretic data that feeds into the construction of an abelian category $\mathcal{A}(G)$ for all toral groups $G$. It makes these explicit for toral subgroups of rank 2 connected groups. It is the first paper in a series of 5 constructing an algebraic category $\mathcal{A}(SU(3))$ and showing it gives an algebraic model for rational $SU(3)$-spectra. This series gives a concrete illustrations of general results in small and accessible examples.
The paper [10] (which does not logically depend on the present paper) constructs algebraic models for all relevant 1-dimensional blocks (which includes all $\mathcal{A}(G)$ 1-dimensional groups $G$). The paper [12] uses the present paper as input to give an algebraic model for the maximal torus normalizer in $U(2)$.
The paper [9] assembles this information and that from [2] to give an abelian category $\mathcal{A}(U(2))$ in 7 blocks and shows it is an algebraic model for rational $U(2)$-spectra. Finally, the paper [8] constructs $\mathcal{A}(SU(3))$ in 18 blocks and shows it is equivalent to the category of rational $SU(3)$-spectra. The most complicated parts of the model for $G=U(2)$ and $G=SU(3)$ are the toral blocks, which are based on the work in the present paper.
This series is part of a more general programme. Future installments will be blocks with Noetherian Balmer spectra [5] and blocks with all Weyl groups finite [11]. The implications of the methods here for the structure of general blocks are explained in [13]. Building on this, an account of the general nature of the models is in preparation [4], and it is hoped that this will be the basis of the proof that the category of rational $G$-spectra has an algebraic model in general.
Part I Subgroups of toral groups (generalities)
The purpose of Part 1 is to describe a method for the classification of conjugacy classes of subgroups of a toral group $G$, which we will illustrate in Part 2 by applying it to all rank 2 toral groups occurring in connected compact Lie groups.
2. Pontrjagin duality for subgroups of toral groups
We set the scene by partitioning subgroups by their image in the component group, and introducing Pontrjagin duality.
2.A. Crude partition
First we observe that for toral groups we can always partition subgroups by their image in the component group. As before we consider a toral group
$1\longrightarrow\mathbb{T}\longrightarrow G\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}W\longrightarrow 1.$
Lemma 2.1.
For a toral group $G$ as above, the space $\mathfrak{X}_{G}=\mathrm{Sub}(G)/G$ of conjugacy classes of subgroups of the toral group $G$ is partitioned into pieces, $\mathcal{V}^{G}_{\overline{H}}$ one for each conjugacy class of subgroups $\overline{H}$ of $W$.
If $\overline{H}\subseteq W$, the set
$\mathcal{V}_{\overline{H}}^{G}=\{(K)\;|\;\pi(K)=\overline{H}\}$
is clopen in the Hausdorff metric topology and closed under passage to cotoral subgroups. Furthermore, $\mathcal{V}^{G}_{\overline{H}}$ is dominated by $\pi^{-1}(\overline{H})$ in the sense that it consists of all subgroups cotoral in $\pi^{-1}(\overline{H})$.
Accordingly, the Balmer spectrum with its Zariski topology is a coproduct
$\mathfrak{X}_{G}=\coprod_{(\overline{H})}\mathcal{V}_{\overline{H}}^{G}.$
Proof : The map $\pi$ induces $\pi_{*}:\mathrm{Sub}(G)\longrightarrow\mathrm{Sub}(W)$, which is continuous in the h-topology and compatible with conjugation.
Since a torus is connected, we see that if $K$ is cotoral in $H$ then $\pi_{*}(K)=\pi_{*}(H)$. Conversely all subgroups $H$ in $\mathcal{V}^{G}_{\overline{H}}$ are cotoral in $\pi^{-1}(\overline{H})$. It is clear that $H$ of $G$ lies inside $\pi^{-1}(\overline{H})$ where $\overline{H}=\pi(H)$. If we pick a (set theoretic) splitting $\sigma$ of $H\longrightarrow\overline{H}$ then $H$ is generated by $S=H\cap\mathbb{T}$ and $\sigma(\overline{H})$ so $H$ is normal in $\pi^{-1}(\overline{H})$ and $\pi^{-1}(\overline{H})/H$ is a quotient of $\mathbb{T}$ and hence a torus. Thus $H\in\mathcal{V}^{G}_{\overline{H}}$.
Thus $\pi_{*}$ is continuous in the Zariski topology. The space $\mathrm{Sub}(W)/W$ is discrete, and so the singletons give a partition into clopen sets. Since $\pi_{*}$ is continuous, their inverse images give a partion of $\mathrm{Sub}(G)/G$. ∎
2.B. Subgroups of a torus
Before considering subtleties, we consider $\mathcal{V}^{G}_{1}$, the subgroups with image the trivial subgroup. More concretely, this is the space of subgroups $S\leq\mathbb{T}$. The group theory is very simple in this case, and the associated part of the Balmer spectrum is always indecomposable. On the other hand, one can argue that the complexity of the space $\mathcal{V}^{G}_{\overline{H}}$ is governed by the rank of the $\overline{H}$-fixed part of the torus $\mathbb{T}^{\overline{H}}$; since $\mathbb{T}$ is of rank $r$, the part of the model over $\mathcal{V}^{G}_{1}$ is often the one of maximal complexity.
It is worth a little extra generality, so we consider an abelian compact Lie group $A$ and apply Pontrjagin duality $B^{*}=\mathrm{Hom}(B,T)$. Associated to a subgroup $B\subseteq A$ we have the subgroup $B^{\dagger}=\ker(A^{*}\longrightarrow B^{*})$ (since we are considering subgroups of a fixed group $A$, the dependence on $A$ is not recorded in the notation).
Lemma 2.2.
The dagger construction is an order reversing isomorphism $\mathrm{Sub}(A)\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\mathrm{Sub}(A^{*})$. The cotoral ordering on $\mathrm{Sub}(A)$ ($C$ is cotoral in $B$ if $B/C$ is a torus) corresponds to the cofree ordering on $\mathrm{Sub}(A^{*})$ ($M$ is cofree in $N$ if $N/M$ is free).
Proof : Pontrjagin duality is an exact functor so the two short exact sequences
$1\longrightarrow B\longrightarrow A\longrightarrow A/B\longrightarrow 1\mbox{ and }1\longleftarrow B^{*}\longleftarrow A^{*}\longleftarrow(A/B)^{*}\longleftarrow 1$
are dual to each other. The group $A/B$ is a torus if and only if $(A/B)^{*}$ is a free abelian group. ∎
Our general strategy is to use Pontrjagin duality and work in terms of sublattices of the dual toral lattice $\Lambda^{0}:=\mathbb{T}^{*}\cong\mathbb{Z}^{r}$. Because $W$ acts on $\mathbb{T}$, it acts on $\Lambda^{0}$, and a $W$-invariant subgroup $S$ of $\mathbb{T}$ corresponds to a $W$-invariant sublattice $\Lambda^{S}:=S^{\dagger}$ of $\Lambda^{0}$. If $S$ is of dimension $d$ then $\Lambda^{S}$ is of corank $d$. We note explicitly that $\Lambda^{S}$ depends on $\mathbb{T}$ as well as $S$, and that $\Lambda^{0}$ is the special case when $S$ is the trivial group.
Remark 2.3.
As we use naturality, it is very important to be clear about variance. We have
$\Lambda^{0}=H^{1}(\mathbb{T};\mathbb{Z})=\mathrm{Hom}(H_{1}(\mathbb{T};\mathbb{Z}),\mathbb{Z})=\mathrm{Hom}(\Lambda_{0},\mathbb{Z}),$
so that the lattices $\Lambda^{0}$ and $\Lambda_{0}$ are dual.
2.C. A little more duality
Both Pontrjagin duality $(\cdot)^{*}=\mathrm{Hom}(\cdot,T)$ and duality of lattices $(\cdot)^{\vee}=\mathrm{Hom}(\cdot,\mathbb{Z})$ play a role in what we do, and we adopt notation to make this clear. We started with a short exact sequence
$1\longrightarrow S\longrightarrow\mathbb{T}\longrightarrow\mathbb{T}/S\longrightarrow 1$
of compact abelian groups, and we want to move as far as possible into the world of free abelian groups. Taking Pontrjagin duals we have
[…]
To restore the variance we take lattice duals to obtain
${((\mathbb{T})^{*})^{\vee}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\Lambda^{0})^{\vee}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\Lambda^{S})^{\vee}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Lambda_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Lambda_{S}}$
In practice we will take $\Lambda_{0}$ to be a fixed copy of $\mathbb{Z}^{r}$ and $\Lambda_{S}$ to be a larger lattice inside $\Lambda_{0}\otimes\mathbb{Q}\cong\mathbb{Q}^{r}$. If we choose bases, and the inclusion $\Lambda^{S}\longrightarrow\Lambda^{0}$ is induced by a matrix $X$ then $\Lambda_{0}\longrightarrow\Lambda_{S}$ is induced by the transposed matrix $X^{t}$ with respect to the dual bases. Note also that we have a diagram
[…]
The point of this last is that in cohomology we have a commutative square
${H^{2}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H^{3}(W;\Lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{2}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H^{3}(W;\Lambda_{S})}$
enabling us to understand the quotient map $\mathbb{T}\longrightarrow\mathbb{T}/S$ in cohomology.
3. Classifying subgroups of toral groups
Considering even the case of $O(2)$ we see that knowing the subgroups is only the first step. For the applications to equivariant cohomology theories, we need a calculational understanding of conjugacy. We give cohomological methods for this, and the examples in Part 2 will provide convincing evidence of their effectiveness.
3.A. Splittings
Let $H$ be a full subgroup of $G$, and note that $S=H\cap\mathbb{T}$ is normal in $H$, so $S$ is a $W$-invariant subgroup of $G$. In particular, we have an extension
$1\longrightarrow S\longrightarrow H\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}W\longrightarrow 1.$
We choose a function $\sigma\colon W\longrightarrow H$ splitting $\pi$; we will say it is normalised if $\sigma(e)=e$. The subgroup $S$ and the section $\sigma$ together specify $H$. Given $G$ and $S$, the condition that a section $\sigma\colon W\to G$ of $\pi\colon G\to W$ determines a subgroup $H$ is that the so called factor set $f_{\sigma}(v,w):=\sigma(vw)^{-1}\sigma(v)\sigma(w)$ takes values in $S$. Indeed, $f_{\sigma}$ is a cycle representing the extension class $\epsilon(H)\in H^{2}(W;S)$.
Conversely, given a subgroup $S$ of $T$ invariant under $W$ and a section $\sigma$ of $\pi$ with factor set taking values in $S$, we write $H(S,\sigma)$ for the associated subgroup of $G$.
Example 3.1.
If $G=O(2)$ then $\mathbb{T}$ is the circle and $W$ is of order 2, and we consider the case that $S$ is cyclic of order $n$. The splittings $\sigma$ correspond to a choice of reflection. We write $\sigma_{\theta}$ for the splitting defined by taking $\sigma_{\theta}(w)$ to be reflection in a line at an angle $\theta$ to the $x$ axis.
The splittings only depend on $\theta$ mod $2\pi$, but $\sigma_{\theta}$ and $\sigma_{\theta+\pi/n}$ specify the same subgroup. Thus the space of subgroups $H$ with $H\cap\mathbb{T}=S$ is the circle specified by $\theta$ mod $\pi/n$. Finally, we observe that if we conjugate $\sigma_{\theta}$ by a rotation by $\phi$ we obtain $\sigma_{\theta+\phi}$ and the circle of subgroups $H$ with $H\cap\mathbb{T}$ of order $n$ constitute a single conjugacy class.
3.B. Subgroups up to conjugacy
We have described the basic invariants of full subgroups of a toral group in terms of the standard form $H(S,\sigma)$, and we have noted that $\sigma$ determines a factor set representing the extension class $\epsilon(H)\in H^{2}(W;S)$. We describe the $G$-conjugacy classes of such subgroups.
We can calculate cohomology with the bar resolution, with cochains $C^{i}(W;M)=\{f:W^{i}\longrightarrow M\}$. In particular, $C^{0}(W;M)=M$ and low dimensional coboundaries are as follows, with $m\in M,g:W\longrightarrow M,f:W^{2}\longrightarrow M$.
$\delta(m)(v)=m^{v}-m,(\delta g)(v,w)=g(v)^{w}-g(vw)+g(w),\\ (\delta f)(u,v,w)=f(u,v)^{w}-f(u,vw)+f(uv,w)-f(v,w).$
Now $\sigma:W\longrightarrow G$ determines $f_{\sigma}(v,w)=\sigma(vw)^{-1}\sigma(v)\sigma(w)$, and since $G$ is a group $\delta f_{\sigma}=0$. We write
$\mathrm{Split}(G|S)=\{\sigma:W\longrightarrow G\;|\;f_{\sigma}(v,w)\in S\}$
for splittings which give subgroups of the form $H(S,\sigma)$.
Lemma 3.2.
(i) If $\sigma:W\longrightarrow G$ is a splitting with $f_{\sigma}(v,w)\in S$ then $H(S,\sigma)$ is a subgroup intersecting $\mathbb{T}$ in $S$.
(ii) $H(S,\sigma)=H(S,\sigma\tau)$ if $\tau\in C^{1}(W;S)$.
(iii) If $g\in C^{1}(W;\mathbb{T})$ then $f_{\sigma g}=f_{\sigma}+\delta g$.
(iv) $tH(S,\sigma)t^{-1}=H(S,\sigma\delta t)$ for $t\in\mathbb{T}=C^{0}(W;\mathbb{T})$.
Proof : Parts (i) and (ii) are clear.
For Part (iii) we calculate
$f_{\sigma g}(v,w)=[\sigma(vw)g(vw)]^{-1}\sigma(v)g(v)\sigma(w)g(w)=\\ g(vw)^{-1}\sigma(vw)\sigma(v)\sigma(w)g(v)^{w}g(w)=f_{\sigma}(v,w)g(vw)^{-1}g(v)^{w}g(w).$
For Part (iv) we calculate
$t\sigma(v)t^{-1}=\sigma(v)\sigma(v)^{-1}t\sigma(v)t^{-1}=\sigma(v)t^{v}t^{-1}.$
Lemma 3.3.
If $\epsilon(G)$ lies in the image of $i^{S}_{*}:H^{2}(W;S)\longrightarrow H^{2}(W;\mathbb{T})$, then $G$-conjugacy classes of subgroups $H(S,\sigma)$ whose extension class lies over $\epsilon(G)$ are in bijection with $H^{1}(W;\mathbb{T}/S)$.
Proof : First, we choose a representative splitting $\sigma_{0}$, with $\delta f_{\sigma_{0}}=0$, giving $f_{\sigma_{0}}]=\epsilon(G)$ in $H^{2}(W;\mathbb{T})$, and since we are assuming $\epsilon(G)$ lifts to $H^{2}(W;S)$, we may assume that $f_{\sigma_{0}}$ takes values in $S$, so that $\sigma_{0}$ gives a subgroup $H=H(S,\sigma_{0})$ and $f_{\sigma_{0}}]=\epsilon(H)$ in $H^{2}(W;S)$.
Any other splitting $\sigma$ takes the form $\sigma=\sigma_{0}\cdot g$ for $g\in C^{1}(W;\mathbb{T})$. Since $f_{\sigma g}=f_{\sigma}+\delta g$ by Lemma 3.2 (iii), any element of this form satisfies $\delta f_{\sigma}=0$ and $f_{\sigma}]=\epsilon(G)$.
The splitting $\sigma$ gives a group over $S$ if and only if $f_{\sigma}$ takes values in $S$, which happens if and only if $\delta g$ takes values in $S$, which is to say $\delta g=0$ in $C^{1}(W;\mathbb{T}/S)$. We also note that by Lemma 3.2 (ii) the subgroup $H(S,\sigma)$ does not change if we multiply by an element $g$ of $C^{1}(W;S)$, so that the subgroup $H$ only depends on $g\in C^{1}(W;\mathbb{T})/C^{1}(W;S)=C^{1}(W;\mathbb{T}/S)$.
To follow the rest of the argument it is useful to refer to the short exact sequence
${C^{0}(W;S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C^{1}(W;S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C^{2}(W;S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C^{0}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C^{1}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C^{2}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C^{0}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C^{1}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C^{2}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$.
of chain complexes.
We have shown that associating $g$ to the subgroup $H(S,\sigma_{0}g)$ gives a bijection
$\{g\in C^{1}(W;\mathbb{T}/S)\;|\;\delta g=0\}=\{H\;|\;H\mbox{ full, }H\cap\mathbb{T}=S\},$
Finally, by By Lemma 3.2 (iv), and using the fact that $C^{0}(W;\mathbb{T})\longrightarrow C^{0}(W;\mathbb{T}/S)$ is surjective, the map induces a bijection
$H^{1}(W;\mathbb{T}/S)=\frac{\{g\in C^{1}(W;\mathbb{T}/S)\;|\;\delta g=0\}}{\{\delta t\;|\;t\in C^{0}(W;\mathbb{T}/S)\}}\cong\{H\;|\;H\mbox{ full, }H\cap\mathbb{T}=S\}/G$
3.C. The ramification sequence
We make frequent use of Lemma 3.3, so it is worth spelling out cohomological interpretations of all stages of the analysis of the ramification exact sequence in cohomology
$H^{1}(W;\mathbb{T})\stackrel{{\scriptstyle a_{S}^{1}}}{{\longrightarrow}}H^{1}(W;\mathbb{T}/S)\longrightarrow H^{2}(W;S)\longrightarrow H^{2}(W;\mathbb{T})\stackrel{{\scriptstyle a_{S}^{2}}}{{\longrightarrow}}H^{2}(W;\mathbb{T}/S).$
The extension class $\epsilon(G)$ lies in the group $H^{2}(W;\mathbb{T})$ and there is a subgroup corresponding to $S$ if and only if $a_{S}^{2}(\epsilon(G))=0$. If so, the conjugacy classes of subgroups are in bijection to $H^{1}(W;\mathbb{T}/S)$, and the isomorphism classes of extensions are in bijection to $\mathrm{cok}(H^{1}(W;\mathbb{T})\longrightarrow H^{1}(W;\mathbb{T}/S))$.
The hierarchy is described by the quotient maps
$\mathrm{Split}(G|S)\longrightarrow\mathrm{Sub}(G|S)\longrightarrow\mathrm{Sub}(G|S)/G\longrightarrow\mathrm{Sub}(G|S)/\mbox{ext-iso}$
Here
$\mathrm{Split}(G|S)=\{\sigma:W\longrightarrow G\;|\;\pi\sigma=1_{W},f_{\sigma}\in C^{2}(W;S),\delta f_{\sigma}=0\}.$
$\mathrm{Sub}(G|S)=\{H\subseteq G\;|\;H\mbox{ full, }H\cap\mathbb{T}=S\}.$
We note that every full subgroup $H$, is preserved by conjugation by an element $\sigma(v)$, so the relation of $G$ conjugacy agrees with $\mathbb{T}$-conjugacy for full subgroups. The advantage here is that since $\mathbb{T}$ is abelian, conjugation by an element of $\mathbb{T}$ gives an isomorphism of extensions.
We may describe the entire process in terms of cohomological data.
If we choose one element $\sigma_{0}\in\mathrm{Split}(G|S)$ then any other element is $\sigma_{0}\cdot g$ where $g\in C^{1}(W;\mathbb{T})$ has $\delta g=0$ in $C^{2}(W;\mathbb{T}/S)$. If $\sigma_{0}$ has the property that $f_{\sigma_{0}}\in C^{2}(W;S)$ then we have a map
$\kappa_{0}:C^{1}(W;\mathbb{T})\longrightarrow\mathrm{Sub}(G|S)/G$
depending on the choice of $\sigma_{0}$ defined by
$\kappa_{0}(g)=\sigma_{0}\cdot g.$
In the following diagram, all vertical maps are quotient maps and we have established the following horizontal isomorphisms
${\mathrm{Split}(G|S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Split}(G|S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Sub}(G|S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Split}(G|S)/C^{1}(W;S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${Z^{1}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\kappa_{0}}$${\mathrm{Sub}(G|S)/G\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Split}(G|S)/C^{1}(W;S),B^{1}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H^{1}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\kappa_{0}}$${\mathrm{Sub}(G|S)/\mbox{ext-iso}}$${\mathrm{Split}(G|S)/C^{1}(W;S),Z^{1}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H^{1}(W;\mathbb{T}/S)/\mathrm{im}H^{1}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\kappa_{0}}$
At the bottom right, using $\sigma_{0}$ again, $H^{1}(W;\mathbb{T}/S)/\mathrm{im}H^{1}(W;\mathbb{T})$ is the subset of $H^{2}(W;S)$ consisting of classes $\epsilon(H)$ which map to $\epsilon(G)\in H^{2}(W;\mathbb{T})$.
3.D. Variation of conjugacy classes
One might naively think that the functor $H^{1}(W;\mathbb{T}/S)$ is continuous in $S$, but this far from true.
Example 3.4.
We observe that $H^{1}(W;\mathbb{T}/S)$ is not a continuous function of $S$. Indeed, we may suppose that $G$ is the 2-torus with the group $W$ of order 2 acting trivially on one factor and as inversion on the other. This means $H_{1}(\mathbb{T})=\mathbb{Z}\oplus\tilde{\mathbb{Z}}$ and we take
$\Lambda_{1}^{n}=\langle(n,0),(0,n)\rangle=n\Lambda^{0},\Lambda_{2}^{n}=\langle(n,n),(n,-n)\rangle.$
Thus we find
$\Lambda^{n}_{2}\supset\Lambda^{2n}_{1}\supset\Lambda^{2n}_{2},$
but $\Lambda^{n}_{1}\cong\mathbb{Z}\oplus\tilde{\mathbb{Z}}$ whilse $\Lambda^{n}_{2}\cong\mathbb{Z}[W]$ and $H^{2}(W;\Lambda^{n}_{1})=\mathbb{Z}/2$ whilst $H^{2}(W;\Lambda^{n}_{2})=0$.
Now take $S^{n}_{i}=\ker(T^{2}\longrightarrow(\Lambda^{n}_{i})^{*})$. We find $S^{n}_{1}\longrightarrow T^{2}$ (with $H^{1}(W;T^{2}/S^{n}_{1})$ being the sequence constant at $\mathbb{Z}/2$) whilst $S^{n}_{2}\longrightarrow T^{2}$, (with $H^{1}(W;T^{2}/S^{n}_{2})$ being the sequence constant at 0).
We must accept the variation and consider a number of ‘sheaves’ over $\mathrm{Sub}(\mathbb{T})$.
  • the tautological sheaf $\gamma$ over $\mathrm{Sub}(\mathbb{T})$ with the group $S$ over the point $S$
  • the constant sheaf $\mathbb{T}$ over $\mathrm{Sub}(\mathbb{T})$
  • the quotient sheaf $\mathbb{T}/\gamma$ over $\mathrm{Sub}(\mathbb{T})$ with the group $\mathbb{T}/S$ over the point $S$
Thus we have a short exact sequence
$0\longrightarrow\gamma\longrightarrow\mathbb{T}\longrightarrow\mathbb{T}/\gamma\longrightarrow 0$
of sheaves of abelian groups. There are associated sheaves of chains, and a long exact sequence of homology sheaves. Of course for any abelian group we have $H^{1}(W;A)=\mathrm{Hom}(W,A),$ so this is rather simple in codegree 1, and $H^{1}(W;\mathbb{T}/S)$ takes only finitely many different values as $S$ varies.
A rather crude summary of the situation may be helpful.
Corollary 3.5.
Up to finite error, conjugacy classes of full subgroups are given by $W$-submodules of the dual toral lattice $\Lambda^{0}=H^{1}(\mathbb{T})$. More precisely the map
$\mathrm{Sub}(G|\mathrm{full})/G\longrightarrow\mbox{$W$-$\mathrm{Sub}(\mathbb{T})$}\cong\mbox{$W$-$\mathrm{Sub}(\mathbb{T}^{*})$}=\mbox{$W$-$\mathrm{Sub}(\Lambda^{0})$}$
taking $H$ to $(H\cap\mathbb{T})^{\dagger}$ is almost surjective (in the sense that it there is an $n$ so that all $S$ containing $\mathbb{T}[n]$ lie in the image) and has finite fibres of bounded cardinality.
Remark 3.6.
Based on this, we depict the space $\mathfrak{X}_{G}$ lying over $W$-$\mathrm{Sub}(\Lambda^{0})$.
Proof : For the almost surjectivity, we use the fact that $H^{2}(W;\mathbb{T})\cong H^{3}(W;\Lambda_{0})$ is finite. The element $\epsilon(G)$ is therefore of finite order and there is an $n$ so that $a^{2}_{S}$ annihilates it whenever $S$ contains $\mathbb{T}[n]$.
We have explained that, when non-empty, the finite fibre over $S$ is the finite group $H^{1}(W;\mathbb{T}/S)\cong H^{2}(W;\Lambda_{S})$. There are usually infinitely many different lattices $\Lambda_{S}$, but up to isomorphism of $\mathbb{Z}W$-modules there are only finitely many. ∎
4. Normalizers of subgroups of a toral group
The essential datum attached to each subgroup is its Weyl group. We describe a systematic approach to calculation in terms of our classification of subgroups.
4.A. Normalizers direct
We describe conjugates and normalizers of subgroups in terms of the standard form.
Lemma 4.1.
Consider a fixed $W$-invariant subgroup $S$.
(i) Two subgroups $H(S,\sigma)$ and $H(S,\sigma^{\prime})$ specify the same subgroup if and only if $\sigma(w)=\sigma^{\prime}(w)\tau(w)$ for some function $\tau\colon W\longrightarrow S$.
(ii) Conjugation by $t\in\mathbb{T}$ is given by
$tH(S,\sigma)t^{-1}=H(S,\sigma\delta(t))$
where $\delta(t)\colon W\longrightarrow T$ is defined by $\delta(t)(v)=t^{v}t^{-1}$.
(iii) The normalizer of a full subgroup $H(S,\sigma)$ is given by
$N_{G}(H(S,\sigma))=H(S^{+},\sigma)\mbox{ where }S^{+}=\{t\;|\;t^{w}t^{-1}\in S\;\forall\;w\in W\}.$
The subset $S^{+}$ of $T$ is a subgroup.
Proof : Parts (i) and (ii) were proved in Lemma 3.2. Part (iii) follows. ∎
We are particularly interested in subgroups with finite index in their normalizer. We note that if we have a containment $H\subseteq H^{\prime}$ of full subgroups, then $S=H\cap T\subseteq H^{\prime}\cap T=S^{\prime}$ and we may choose the same section $\sigma$ so the containment is of the form $H=H(S,\sigma)\subseteq H(S^{\prime},\sigma)=H^{\prime}$. Such a containment occurs if and only if $S\subseteq S^{\prime}$ and that in that case
[…]
Indeed, $H(S,\sigma)$ is normal in $H(S^{\prime},\sigma)$ if and only if $H(S^{\prime},\sigma)\subseteq H(S^{+},\sigma)$ by Lemma 4.1, which is to say $S^{\prime}\subseteq S^{+}$.
Proposition 4.2.
The subgroup $H(S,\sigma)$ is of finite index in its normalizer if and only if $\dim(S^{W})=\dim(T^{W})$.
Proof : By Lemma 4.1, we have $N_{G}(H(S,\sigma))=H(S^{+},\sigma)$, so taking $S^{\prime}=S^{+}$ in (1), we see that $H(S,\sigma)$ is of finite index in its normalizer if and only if $S^{+}/S$ is finite, or equivalently if $LS^{+}=LS$. From the description of $S^{+}$ in Lemma 4.1, we get
$LS^{+}=\{x\in LT\;|\;x-wx\in LS\mbox{ for all }w\in W\}.$
Since $S$ is a $W$-invariant subgroup of $T$ we see
$LS=\bigoplus_{\alpha}U_{\alpha}\mbox{ with }U_{\alpha}\leqslant V_{\alpha}.$
Picking a submodule $U_{\alpha}^{\prime}\leqslant V_{\alpha}$ so that $V_{\alpha}=U_{\alpha}\oplus U^{\prime}_{\alpha}$, this implies that if $x\in LS^{+}$ has components
$(x_{\alpha},x_{\alpha}^{\prime})\in U_{\alpha}\oplus U_{\alpha}^{\prime}=V_{\alpha}$
then the requirement that $LS^{+}=LS$ gives $x_{\alpha}^{\prime}-wx_{\alpha}^{\prime}=0$ for all $w\in W$ and all $\alpha$. Thus $x_{\alpha}^{\prime}=0$ for $\alpha\neq 1$ and
$LS^{+}/LS=V_{1}/U_{1}.$
This in turn is the case if and only if $\dim(S^{W})=\dim(T^{W})$. ∎
4.B. Normalizers by Pontrjagin duality
We describe how to calculate the normalizer of a subgroup of a toral group from the lattice data.
As usual we suppose $H=H(S,\sigma)$ with $S\subseteq\mathbb{T}$. We have seen that $N_{G}(H(S,\sigma))=H(S^{+},\sigma)$ where $S^{+}=\{t\in\mathbb{T}\;|\;t^{-1}t^{v}\in S\mbox{ for all }v\in W\}$.
In the dual setting, we may suppose $\Lambda^{S}=S^{\dagger}$ and $\Lambda^{S}_{+}:=(S^{+})^{\dagger}$ where $\Lambda^{S}_{+}\subseteq\Lambda^{S}$. The calculation of $\Lambda^{S}_{+}$ is a matter of linear algebra, as we now explain.
We suppose $\mathbb{T}$ is a torus of rank $r$ and start with the case that $W$ is cyclic generated by $g$. First of all, we choose bases so that $\mathbb{T}^{*}=\mathbb{Z}^{r}$, and we identify $\mathbb{T}$ with its double dual $\mathbb{T}\cong(\mathbb{Z}^{r})^{*}$. We will write $\phi:\mathbb{Z}^{r}\longrightarrow T$ for an element of $S$ and name the images of the standard basis $\phi(e_{i})=z_{i}\in T$ for $1\leq i\leq r$. The last piece of notation is that in this basis $g$ is given by the matrix $A(g)$.
Now suppose $S$ is of dimension rkr-k and $\Lambda^{S}$ has basis $\lambda^{1},\ldots,\lambda^{k}$, and we express the $\lambda^{i}$ in terms of the standard basis: $\lambda^{i}=(x^{i}_{1},\ldots,x^{i}_{r})$. We write $L$ for the matrix with columns $\lambda^{1},\ldots,\lambda^{k}$. The condition that $\phi$ lies in $S$ is that $\phi(\lambda^{i})=1$ for $1\leq i\leq k$, which is to say
$w_{1}^{x^{i}_{1}}w_{2}^{x^{i}_{2}}\cdots w_{r}^{x^{i}_{r}}=1\mbox{ for }i=1,\ldots,k.$
In matrix notation,
$(w_{1},\ldots,w_{r})L=(1,\ldots,1).$
Converting to additive notation, if $w_{j}=e^{2\pi is_{j}}$ for $1\leq j\leq r$, we may rewrite this in a purely additive setting:
$(s_{1},\ldots,s_{r})L\in\mathbb{Z}^{2}.$
Moving on to $S_{+}$, the condition that $\theta$ lies in $S_{+}$ is that $\phi:=\theta^{-1}\theta^{g}$ lies in $S$. Including variables for the functions on $\mathbb{Z}^{r}$ this states that $\phi(x_{1},\ldots,x_{r})=\theta A(g)(x_{1},\ldots,x_{r})/\theta(x_{1},\ldots,x_{r})$ lies in $S$, which we may check by evaluating on the generators of $\Lambda^{S}$: we require
$\theta A(g)\lambda^{i}=\theta\lambda^{i}\mbox{ in $T$ for }1\leq i\leq k$
In additive matrix form this states that the product
$(s_{1},\ldots,s_{r})M(g)L$
is an integer vector where M(g)=A(g)IM(g)=A(g)-I. In other words $S_{+}$ is dual to the lattice $M(g)L$. If $S$ is finite (so that $L$ is a non-singular square matrix) the order of $S^{+}/S$ (if finite) is $|\det(M(g))|$, but the structure will depend on the lattice and the actual matrix.
If $W$ is not cyclic, we require that the condition holds each generator $g$ of $W$, so we find $S_{+}$ is dual to the lattice $\Lambda^{S}_{+}=\sum_{g}M(g)L$.
Example 4.3.
(i) In rank 1, if $W=\langle g\rangle$ is of order 2 and $g$ acts as 1-1 and $S$ is cyclic of order $m$ then $\Lambda^{S}=m\mathbb{Z}$ and the condition for membership of $S$ is
$rm\in\mathbb{Z}$
so that $r$ is a multiple of $1/m$. The condition for membership $S^{+}$ is
$r(-2)m\in\mathbb{Z}$
so that $r$ is a multiple of $1/(2m)$ and $S^{+}$ is cyclic of order $2m$.
(ii) In rank 2, if $W=\langle g\rangle$ is of order 2 and $T^{*}=\mathbb{Z}\oplus\mathbb{Z}$, then $g$ acts as the identity and the matrix $M=0$, so $S^{+}=T^{2}$ for every $S$.
(iii) If $W=\langle g\rangle$ is of order 2 and $T^{*}=\mathbb{Z}\oplus\tilde{\mathbb{Z}}$ then
$M:=\left(\begin{array}[]{cc}0&0\\ 0&-2\end{array}\right).$
The Weyl group of each finite $S$ has identity component $T$, but its structure depends on the lattice (see Section 8).
(iv) If $W=\langle g\rangle$ is of order 2 and $T^{*}=\tilde{\mathbb{Z}}\oplus\tilde{\mathbb{Z}}$, then
$M:=\left(\begin{array}[]{cc}-2&0\\ 0&-2\end{array}\right)$
and the Weyl group of any finite $S$ is $C_{2}\times C_{2}$.
(v) If $W=\langle g\rangle$ is of order 3 and $(T^{2})^{*}=\mathbb{Z}[\omega]$ $g$ acts via $\left(\begin{array}[]{cc}0&-1\\ 1&-1\end{array}\right)$ giving
$M:=\left(\begin{array}[]{cc}-1&-1\\ 1&-2\end{array}\right)$
and the Weyl group of every finite $S$ is of order 3.
5. The topology on spaces of subgroups
One expects that the structure on $\mathrm{Sub}(G)/G$ relevant to the building of the model is the Zariski topology. As a general fact, spectral spaces can be described using the constructible topology and the specialization order; this is described in [1] where it is also shown that the constructible topology is the $h$-topology (induced by the Hausdorff metric) and the specialization order is the cotoral order. However, for this spectral space [7] shows it is also determined by the cotoral ordering together with the $h$-topology on the spaces $\Phi H$ ($H$-conjugacy classes of subgroups of $H$ with finite $H$-Weyl group). To see the advantage of the second point of view, we note that for Noetherian spaces (finite central extensions of a torus) it means the topology is determined by the poset of subgroups under the cotoral ordering.
In any case, Section 4 determined the cotoral ordering, and in this section we consider the Hausdorff metric topology on the space $\mathcal{V}^{G}_{\mathrm{full}}$ of conjugacy classes of full subgroups.
5.A. Reductions
Since we are working with the Hausdorff metric topology it is natural to describe the topology on the space of subgroups in terms of convergent sequences, and we will do this in terms of the associated lattices.
First, we observe that we may restrict attention to full subgroups.
Lemma 5.1.
If the sequence $(H_{n})$ of subgroups of $G$ is convergent to a full subgroup $H_{*}$, then almost all the subgroups $H_{n}$ are full (and hence $H_{n}=H(S_{n},\sigma_{n})$).
Proof : Clear since the projection $G\longrightarrow W$ is continuous, and the topology on $W$ is discrete. ∎
Next, we may reduce to consideration of the subgroups of the torus.
Lemma 5.2.
If we have a sequence $H_{n}=H(S_{n},\sigma_{n})$ of full subgroups of $G$, then $H_{n}\longrightarrow H_{*}$ if and only if $S_{n}\longrightarrow S_{*}:=H_{*}\cap\mathbb{T}$.
Proof : The map taking intersections with $\mathbb{T}$ is continuous, so if $H_{n}\longrightarrow H$ then $S_{n}\longrightarrow S$.
Now suppose $S_{n}\longrightarrow S$. By Montgomery-Zippin, after conjugation and omitting finitely many terms, we may assume $H_{n}\subseteq H$. This means $H=S\cdot H_{n}$, and $d(\sigma_{n}(v)\cdot S_{n},\sigma(v)_{n}\cdot S)=d(S_{n},S)$. It follows that $d(H_{n},H)\leq d(S_{n},S)$. ∎
Finally, we may reduce to consideration of sequences tending to a torus.
Lemma 5.3.
(i) If $S\subseteq\mathbb{T}$ then there is a neighbourhood of $S$ consisting of subgroups $S^{\prime}$ with $S^{\prime}$ cotoral in $S$.
(ii) If $S^{\prime}$ is cotoral in $S$ then $d(S,S^{\prime})=d(S_{e},S_{e}\cap S^{\prime})$.
Proof : (i) Since the group is abelian, by Montgomery-Zippin, there is a neighbourhood of $S$ consisting of subgroups of $S$. The different components are separated by some distance $d$, and any subgroup $S^{\prime}$ closer to $S$ than this must contain points in each component of $S$.
(ii) The invariance of translation shows that the distance between $S_{e}$ and $S_{e}\cap S^{\prime}$ is the same as that between $tS_{e}$ and $tS_{e}\cap S^{\prime}$. ∎
5.B. Tori
It remains to understand in terms of lattices when subgroups of a torus converge to the torus. For a non-zero lattice $\Lambda\subseteq\mathbb{Z}^{r}$ we write
$\mu(\Lambda):=\min\{|\lambda|_{\infty}|0\neq\lambda\in\Lambda\}$
for the shortest vector in the max metric.
Proposition 5.4.
For a sequence $S_{n}$ of subgroups of $\mathbb{T}$,
$S_{n}\longrightarrow\mathbb{T}\Leftarrow\Rightarrow\mu(\Lambda^{S_{n}})\longrightarrow\infty.$
In fact we only need to estimate the distance between a subgroup $S_{*}$ and a codimension 1 subgroup $S$ of $\mathbb{T}$. One should think of the a circle of slope $p/q$ (in lowest terms) in a 2-torus will be within about $1/2q$ of every other point. For a codimension 1 subgroup, $\Lambda^{S}=\langle\theta\rangle$ where $\theta=(a_{1},\ldots,a_{r})\neq 0$, we have $\mu(S)=\max\{|a_{i}|\;|\;0\leq i\leq r\}.$
Lemma 5.5.
There are numbers $0<a<b$ (depending only on $r$) so that if $S\subseteq\mathbb{T}$
$\frac{a}{\mu(S)}<d(S,\mathbb{T})<\frac{b}{\mu(S)}$
Proof : The condition for $\theta\in\Lambda^{S}$ is that $\theta(\lambda)=1$ for $\lambda\in\Lambda^{S}$.
For notational simplicity suppose that $\lambda=(a_{1},a_{2},\ldots,a_{r})$ and $a_{1}>0$ is the entry of the largest modulus. Then if the point $P_{b}:=(x_{1}+b,x_{2},\ldots,x_{r})$ lies in $\ker(\theta)$ for $b=0$, then the point $P_{n/a_{1}}$ lies in $\ker(\theta)$ for every integer $n$. This shows that the distance from a point to an element of $\ker(\theta)$ is $\leq 1/(2a_{1})$. A little trigonometry shows the maximum distance is $\geq 1/(2\sqrt{r}a_{1})$. ∎
This is sufficient to give an estimate for arbitrary subgroups, and we may now prove the proposition.
Proof : First of all, if $\theta\in\Lambda^{S}$ then $S\subseteq\ker(\theta)$ so that
$d(S,\mathbb{T})\geq d(\ker(\theta),\mathbb{T})>a/\mu(\theta)\geq a/\mu(\Lambda^{S}),$
so that any convergent sequence must have $\mu(\Lambda^{S_{n}})\longrightarrow\infty$.
On the other hand, we may express $S$ as an intersection of codimension 1 subgroups so that
$S=S^{c}\subset S^{c-1}\subset\cdots\subset S^{0}=\mathbb{T}$
and
$d(S,\mathbb{T})\leq\sum_{i=0}^{c-1}d(S^{i},S^{i+1}).$
If $S^{i}=S^{i-1}\cap\ker(\theta_{i})$ then $d(S^{i},S^{i-1})<d(\ker(\theta_{i}),\mathbb{T})<b/\mu(\theta_{i})$, so that if $\mu(\Lambda^{S_{n}})\longrightarrow\infty$ then $d(S_{n},\mathbb{T})\longrightarrow 0$. ∎
Corollary 5.6.
For any subgroup $S\subseteq\mathbb{T}$ there is a base of neighbourhoods $\mathcal{N}_{t}(S)$ of $S$ so that the corresponding collection of lattices is
$\mathcal{N}_{t}(\Lambda^{S})=\{\Lambda^{\prime}\supseteq\Lambda^{S}\;|\;\mu(\Lambda^{\prime})>1/t\}.$
Remark 5.7.
Since the number of vectors with all terms <N<N is finite for any $N$, this base of neighbourhoods generates the same topology as the cofinite base of neighbourhoods so the topology is entirely determined by containment of subgroups.
This lets us describe the space $\mathrm{Sub}(\mathbb{T})$ as a compactification of the discrete space $\mathfrak{X}_{0}$ of finite subgroups. In fact we let
$\mathcal{F}=\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\cdots\subseteq\mathcal{F}_{r}=\mathrm{Sub}(\mathbb{T})$
denote the dimension filtration, with
$\mathcal{F}_{i}=\{S\subseteq\mathbb{T}\;|\;\dim(S)\leq i\}.$
Now we define $\mathfrak{X}_{i}$ to be the set $\mathcal{F}_{i}$ with a suitable topology. Indeed, we take $\mathfrak{X}_{0}:=\mathcal{F}_{0}$ to be discrete and we define $\mathfrak{X}_{i}\longrightarrow\mathfrak{X}_{i+1}$ to be a partial compactification. Indeed, $\mathfrak{X}_{i+1}$ is obtained from $\mathfrak{X}_{i}$ by adding the subgroups of dimension $i+1$, and we give it a locally cofinite topology: the neighbourhoods of the $(i+1)$-dimensional subgroup $H$ correspond to finite sets $\tau$ of subgroups of $H$
$U_{\tau}:=\{K\;|\;H\supseteq K\not\in\tau\}.$
One sees immediately that any set of subgroups with finitely many maximal elements is compact, so the space $\mathfrak{X}=\mathfrak{X}_{r}$ is compact.
Part II Subgroups of toral groups (rank 2 examples)
6. Method and Summary
We describe our general procedure for an explicit determination of space of conjugacy classes of subgroups of a toral group $G$, together with data about the topology, cotoral inclusions and Weyl groups. This gives the Balmer spectrum of finite rational $G$-spectra for a toral group $G$.
When we have described the method, we will give a detailed list of the examples to which we apply it.
6.A. The method
The method may be summarized as follows.
  1. (1)
    Determine the dual toral lattice $\Lambda^{0}$ (as a $W$-module).
  2. (2)
    Classify the $W$-invariant sublattices $\Lambda^{S}$ of $\Lambda^{0}$, and identify their $W$-module structures.
  3. (3)
    Determine the maps $\Lambda^{S}\longrightarrow\Lambda^{0}$ and their duals $k_{S}:\Lambda_{0}\longrightarrow\Lambda_{S}$
  4. (4)
    Calculate $H^{i}(W;\Lambda_{S})$ for $i=1,2$, and the maps induced by $k_{S}$.
  5. (5)
    Feed all this information into the Ramification Exact Sequence of Subsection 3.C:
    ${H^{1}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{a_{S}^{1}}\scriptstyle{\cong}$${H^{1}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H^{2}(W;S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{2}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{a_{S}^{2}}\scriptstyle{\cong}$${H^{2}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H^{2}(W;\Lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{k^{S}_{*}}$${H^{2}(W;\Lambda_{S})}$${H^{3}(W;\Lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{k^{S}_{*}}$${H^{3}(W;\Lambda_{S})}$
    For each $S$ this gives a classification of the $G$-conjugacy classes occurring amongst $H(S,\sigma)$. If $\epsilon(G)$ lifts to $S$ then the conjugacy classes are in bijection with $H^{2}(W;\Lambda_{S})$, partitioned into extension classes by cosets of the image of $H^{2}(W;\Lambda_{0})$.
  6. (6)
    Determine how the cohomological classification of conjugacy classes of subgroups varies with $S$.
  7. (7)
    Display the lattice of inclusions amongst the $W$-invariant lattices $\Lambda^{S}$, with particular attention to those that are cofree, and those which are cofree and $W$-cotrivial.
  8. (8)
    Display the lattice of inclusions amongst the $W$-invariant subgroups $S$ with those that are cotoral, and those which correpond to cotoral inclusions of subgroups $H$.
  9. (9)
    The actual data is the finite cover with $H^{2}(W;\Lambda_{S})$ over $S$: add this data to the diagram.
  10. (10)
    The topology is determined by the subgroups $S$: add any necessary commentary.
  11. (11)
    The Weyl groups are determined by the linear algebra of $\Lambda^{S}$; decorate the diagram accordingly.
  12. (12)
    (Optional) Determine the group structure of the groups $H$. (This can be helpful in determining fusion).
6.B. Catalogue of examples
In the following chart, the first column identifies the section of this paper treating each case. The group $G$ is determined up to extension by the first two entries $(W,\Lambda_{0})$, and we note that in all these cases $\Lambda^{0}\cong\Lambda_{0}$ as $\mathbb{Z}W$-modules. When $(W,\Lambda_{0})$ is the Weyl group and root lattice, this name is used. Otherwise we write $\mathrm{Cub}$ for the standard lattice $\mathbb{Z}^{2}$ aligned so $(1,0)$ is on the $x$-axis and $(0,1)$ is on the $y$-axis, and $FCC$ for the face centred cubic lattice obtained from $\mathrm{Cub}$ by adding a second copy $FCC=\mathrm{Cub}\cup(\mathrm{Cub}+(1/2,1/2))$. Where necessary there is a subscript indicating how $W$ acts.
We remind the reader that everything is encoded in terms of the Ramification Exact Sequence:
$A\stackrel{{\scriptstyle\alpha}}{{\longrightarrow}}B\longrightarrow C\longrightarrow D\stackrel{{\scriptstyle\beta}}{{\longrightarrow}}E.$
The extension class $\epsilon(G)$ lies in the group $D$ and there is a subgroup corresponding to $S$ if and only if $\beta(\epsilon(G))=0$. If so, the conjugacy classes of subgroups are in bijection to $B$, and the isomorphism classes of extensions are in bijection to $\mathrm{cok}(A\longrightarrow B)$. In the 5 rows marked by an asterisk $\alpha$ and $\beta$ are determined by the groups $A,B,D,E$ alone. In the remaining 12 rows, $\alpha$ and $\beta$ may depend on the subgroup $S$ (and not just $\Lambda^{S}$ as a representation of $W$). We recommend that the reader keeps this table to hand when examining each case.
The cohomology calculations in the table will be justified in Section 7 below.
The groups are all elementary abelian, so $2^{4}$ abbreviates $(C_{2})^{4}$.
[…]
7. Cohomology calculations
This section is taken up with group cohomology calculations that feed in to the classification of groups and subgroups: in effect this means we need to calculate $H^{i}(W;\mathbb{T})\cong H^{i+1}(W;\Lambda_{0})$ and $H^{i}(W;\mathbb{T}/S)\cong H^{i+1}(W;\Lambda_{S})$ and for $i=1,2$ for all $W$-invariant sublattices $\Lambda^{S}\subseteq\Lambda^{0}$. (In all our cases the modules $\Lambda^{S}$ are self-dual, so the transition to $\Lambda_{S}$ is immediate).
7.A. Classifying extensions
We will discuss small examples of two types. First we consider the examples of direct interest (i.e., the rank is $\leq 2$ and $W$ is a subgroup of the Weyl group of a connected rank 2 group). We also add a few other groups for perspective.
The basic inputs are the following well known calculations
$H^{*}(C_{n};\mathbb{Z})=\mathbb{Z}[x]/(nx),|x|=2\mbox{ and }H^{*}(\Sigma_{3};\mathbb{Z})=\mathbb{Z}[y]/(3y),|y|=4.$
The other fact we use is that $H^{*}(W;\mathbb{Z}[W])=H^{0}(W;\mathbb{Z}[W])=\mathbb{Z}$ for any finite group $W$.
7.B. Rank 1
We first suppose that $r=1$. Since $GL_{1}(\mathbb{Z})$ is of order 2, we either have a trivial action or a surjection $W\longrightarrow GL_{1}(\mathbb{Z})$.
If the action is trivial then the split extension is a product, $G=T\times W$, and the non-split extensions correspond to $H^{2}(W;T)\cong H^{3}(W;\mathbb{Z})$. If $W$ is a subgroup of $\Sigma_{3}$ these groups are zero so all extensions are split.
If the action is non-trivial, the action is given by a surjective map $W\longrightarrow\overline{W}=GL_{1}(\mathbb{Z})$. If $W$ is of order 2, then the split extension is $G=O(2)$ and there is a unique non-split extension $G=Pin(2)$, and in general the group is obtained from these cases by pullback along $W\longrightarrow\overline{W}$.
7.C. Rank 2
We now suppose $r=2$. We break the discussion according to the component group $W$.
7.C.1. $W$ of order 2
Starting with $W$ of order 2, the indecomposable representations of the group of order 2 are $\mathbb{Z}$ (trivial), $\tilde{\mathbb{Z}}$ (generator acts as 1-1) and the two dimensional representation $\mathbb{Z}W$ [3, 74.3]. We have $H^{3}(W;\mathbb{Z})=H^{3}(W;\mathbb{Z}W)=0$, so extensions if $\Lambda_{0}=\mathbb{Z}^{2}$ or $\mathbb{Z}W$ are split and $G=T^{2}\times W$ in the first case and $T^{2}\rtimes W$ in the second (with $W$ exchanging the two factors). On the other hand, $H^{3}(W;\tilde{\mathbb{Z}})=\mathbb{Z}/2$, where the non-split extension is $Pin(2)$. If $\Lambda_{0}=\mathbb{Z}\oplus\tilde{\mathbb{Z}}$ this gives $G\cong O(2)\times T$ in the split case and $G=Pin(2)\times T$ in the non-split case. If $\Lambda_{0}=\tilde{\mathbb{Z}}\oplus\tilde{\mathbb{Z}}$ there are three non-split cases depending on which element of $T^{2}[2]$is a square.
7.C.2. $W=C_{2}\times C_{2}$
We consider only the toral lattices occuring in rank 2 connected groups. There are two cases to consider.
Case 1: $\Lambda_{0}$ is the $A_{1}\times A_{1}$ lattice, so that $\Lambda_{0}=\mathbb{Z}^{2}=\mathrm{Cub}$ is the standard lattice where the coordinates are $(x,y)$, and $W$ acts by reflection in the axes. We therefore find $\Lambda_{0}=\tilde{\mathbb{Z}}_{x}\oplus\tilde{\mathbb{Z}}_{y}$ where $\tilde{\mathbb{Z}}_{x}$ is the $x$-axis and $\tilde{\mathbb{Z}}_{y}$ is the $y$-axis. Factoring out the reflection in the $x$-axis or the $y$-axis gives two Serre spectral sequences. Comparing the two we see the differentials are zero in degrees affecting $H^{i}(W;\mathbb{Z}_{x})$ with $i\leq 3$ and hence $H^{2}(W;\tilde{\mathbb{Z}}_{x})\cong\mathbb{Z}/2$ and $H^{3}(W;\tilde{\mathbb{Z}}_{x})\cong(\mathbb{Z}/2)^{2}$. This gives the tabulated value. We obtain examples which are products of $SU(2)$ and $SO(3)$; the large number arise since there is a choice about which element of $W$ lifts to something whose square is non-trivial in each factor.
[…]
7.C.3. $W=C_{4}$
[…]
7.C.4. $W=D_{8}$
Here we restrict attention to the case when $\Lambda_{0}$ is the $B_{2}$ lattice, with the Weyl group acting in the standard way. This is arranged so that the groups $C_{4}$ acts as in the previous example and the two copies of the Klein 4-group act in the two ways detailed above.
Considering the extension $1\longrightarrow\langle a,b\rangle\longrightarrow D_{8}\longrightarrow C_{2}\longrightarrow 1$ we see that $H^{i}(D_{8};\Lambda_{0})$ is $022X$ where $X$ is either of order 4 or 8.
Considering the extension $1\longrightarrow C_{4}\longrightarrow D_{8}\longrightarrow C_{2}\longrightarrow 1$, we see that $H^{i}(D_{8};\Lambda_{0})$ is $022X$ where $X$ is of order 2 or 4. Combining these we see that $X$ is of order 4.
7.C.5. $W$ of order 3
If $W=\langle g\rangle$ has order 3, the non-trivial indecomposable representations are $\mathbb{Z}W/(1+g+g^{2})\cong\mathbb{Z}[\omega]$ (where $\omega$ is a cube root of $1$), and the three dimensional representation $\mathbb{Z}W$ [3, 74.3]. We record the cohomology in all cases.
The cohomology calculations are as follows.
Lemma 7.1.
  1. (1)
    The cohomology ring is $H^{*}(W;\mathbb{Z})=\mathbb{Z}[x]/(3x)$ where $|x|=2$,
  2. (2)
    The cohomology of $\mathbb{Z}[\omega]$ is $H^{*}(W;\mathbb{Z}[\omega])=\Sigma\mathbb{Z}/3[x]$.
We conclude that extensions corresponding to $\Lambda_{0}=\mathbb{Z},\mathbb{Z}W$ all split, whilst for $\Lambda_{0}=\mathbb{Z}[\omega]$ there is a split extension and two isomorphic non-split extensions.
Proof : The statements for the modules $\mathbb{Z}$ and $\mathbb{Z}W$ are well known. The statement for $\mathbb{Z}[\omega]$ is immediate from the short exact sequence
$0\longrightarrow\mathbb{Z}\longrightarrow\mathbb{Z}W\longrightarrow\mathbb{Z}[\omega]\longrightarrow 0.$
Thus with $W$ of order 3 and $\mathbb{T}$ of rank $\leq 3$, $G=\mathbb{T}\rtimes W$, unless $G$ has a summand $\mathbb{Z}[\omega]$ where there are two isomorphic non-split extensions.
7.C.6. $W=D_{6}$
The one dimensional representations are again $\mathbb{Z}$ or $\tilde{\mathbb{Z}}$. If $\Lambda$ is of rank 2 then either $C_{3}$ acts trivially (and $\Lambda$ is inflated from the group $\Sigma_{3}/C_{3}$ of order 2) or else $M$ is a copy of $\mathbb{Z}[\omega]$ as a $C_{3}$-module. The reflection $\tau$ in the $x$-axis cannot act as a scalar (since $\tau g\tau=g^{-1}$), so it must have eigenvalues +1+1 and 1-1. The first example is when $\tau$ acts as complex conjugation on $\mathbb{Z}[\omega]$.
Lemma 7.2.
There are two $D_{6}$-modules with underlying $C_{3}$-module $\mathbb{Z}[\omega]$ the first $\mathbb{Z}[\omega]^{\prime}$ in which $\tau$ acts as complex conjugation (reflection in the line through $v=1$), and the second $\mathbb{Z}[\omega]^{\prime\prime}$ in which $\tau$ acts as reflection in the line through $v=2+\omega$.
The one to occur in the $A_{2}$ case is $\mathbb{Z}[\omega]^{\prime}$.
Proof : The rational representation $\Lambda\otimes\mathbb{Q}$ is a sum of $\mathbb{Q}$ and $\tilde{\mathbb{Q}}$, so $\Lambda^{\tau}$ is of rank 1. If $v\in\Lambda$ is fixed by $\tau$ then $\omega^{2}v=\tau g\tau v=\tau\omega v$ meaning that $\tau$ acts as reflection in the line through $v$. Up to isomorphism we may assume the argument of $v$ lies between 0 and $\pi/3$. The only angles taking $1$ to another lattice point are $v=1$ or $v=2+\omega$. ∎
The cohomology calculations follow from Lemma 7.1.
Lemma 7.3.
We have $H^{*}(\Sigma_{3};M)=H^{*}(C_{3};M)^{C_{2}}$. This gives the following values.
  1. (1)
    ($M=\mathbb{Z}$) $H^{*}(D_{6};\mathbb{Z})=\mathbb{Z}[y]/(3y)$ with $|y|=4,y=x^{2}$.
  2. (2)
    ($M=\mathbb{Z}$) $H^{*}(D_{6};\tilde{\mathbb{Z}})=\Sigma_{2}\mathbb{Z}/3[y]$ with $|y|=4$.
  3. (3)
    ($M=\mathbb{Z}[\omega]^{\prime}$) $H^{*}(D_{6};\mathbb{Z}[\omega]^{\prime})=\Sigma_{1}\mathbb{Z}/3[y]$ with $|y|=4$.
  4. (4)
    ($M=\mathbb{Z}[\omega]^{\prime\prime}$) $H^{*}(D_{6};\mathbb{Z}[\omega]^{\prime\prime})=\Sigma_{3}\mathbb{Z}/3[y]$ with $|y|=4$.
Proof : For $\mathbb{Z}[\omega]^{\prime},\mathbb{Z}[\omega]^{\prime\prime}$ we see that $H^{1}(C_{3};\mathbb{Z}[\omega])=\mathbb{Z}[\omega]/(1-\omega)$ is generated by the image of $1$. This is fixed if $v=1$ and it is negated if $v=2+\omega$. The generator in degree 3 is obtained by multiplying with a generator of $H^{2}(C_{3};\mathbb{Z})$, which is negated. ∎
7.C.7. $W=D_{12}$
In this case we restrict attention to the case when $\Lambda_{0}$ is the $G_{2}$ lattice, with the Weyl group acting in the standard way. This is arranged so that the group $D_{6}$ acts as in $A_{2}$: in previous notation this is $\Lambda_{0}|_{D_{6}}=\mathbb{Z}[\omega]^{\prime}$. Hence
$H^{i}(D_{12};\Lambda_{0})=H^{i}(D_{6};\mathbb{Z}[\omega]^{\prime})^{C_{2}}.$
This shows it vanishes for $i=2,3$ as tabulated.
7.D. $W$ of prime order
We insert one more generic answer to hint at the complexities in higher rank. By [3, 74.3], the indecomposable representations of a group of prime order $p$ are $\mathbb{Z}$, $\mathfrak{a}$ for a fractional ideal $\mathfrak{a}$ in $\mathbb{Q}W$ (of rank p1p-1) or $\Lambda_{\mathfrak{a}}$ of rank $p$, where
$0\longrightarrow\mathfrak{a}\longrightarrow\Lambda_{\mathfrak{a}}\longrightarrow\mathbb{Z}\longrightarrow 0;$
if $\Lambda_{\mathfrak{a}}=\mathfrak{a}\oplus\mathbb{Z}\cdot y$ additively, we may suppose $gy=1+y$. This means that $\mathbb{T}$ is a product of tori of dimension 1,p11,p-1 and $p$.
8. The case $W=C_{2}$, $\Lambda_{0}=\mathbb{Z}\oplus\tilde{\mathbb{Z}}$
This covers the groups $T\times O(2)$ and $T\times Pin(2)$.
8.A. Preamble
Suppose that $W$ is of order 2 and the toral lattice $\Lambda_{0}=\mathbb{Z}\oplus\tilde{\mathbb{Z}}$. In this case, we have $H^{2}(W;T^{2})=H^{3}(W;\Lambda_{0})=\mathbb{Z}/2$. Furthermore, the torus itself splits following the positive and negative eigenspaces so $T^{2}=Z\times\tilde{T}$, where $Z$ is the identity component of the centre. The split and non-split examples are the groups $T\times O(2)$ and $T\times Pin(2)$.
8.B. Lattices and Weyl groups
For subgroups we must classify $W$-invariant lattices in $\mathbb{Z}\oplus\tilde{\mathbb{Z}}$ and identify their Weyl groups. We have described the method in Subsection 4.B above. For this lattice the matrix $M$ is given by
$M=\left(\begin{array}[]{cc}0&0\\ 0&-2\end{array}\right)$
Lemma 8.1.
(i) The unique rank 0 sublattice of $\Lambda^{0}$ is 0, and it corresponds to $S=T^{2}$. It has trivial Weyl group.
(ii) The rank 1 lattices are of two types
(a) The lattice $\Lambda_{+}(m)$ with basis $\lambda=(m,0)$ for an integer $m\geq 1$. As a $W$-module this is $\mathbb{Z}$. It corresponds to $S=C_{m}\times\tilde{T}$ which has Weyl group the circle $Z/C_{m}$.
(b) The lattice $\Lambda_{-}(n)$ with basis $\lambda=(0,n)$ for an integer $n\geq 1$. As a $W$-module this is $\tilde{\mathbb{Z}}$. It corresponds to $S=Z\times C_{n}$ and has Weyl group $D_{4n}/D_{2n}$ of order 2.
Proof : The only comment to make is that if $\Lambda$ is an invariant rank 1 lattice then $g\Lambda=\Lambda$ so that a generator $\lambda$ is taken to $\pm\lambda$. For the Weyl groups we note $M(m,0)^{t}=(0,0)^{t}$ but $M(0,n)^{t}=(0,-2n)^{t}$. ∎
Lemma 8.2.
There are two families of rank 2 $W$-invariant lattices in $\mathbb{Z}\oplus\tilde{\mathbb{Z}}$.
Type 1: $\Lambda_{1}(m,n)=m\mathbb{Z}\oplus n\tilde{\mathbb{Z}}$ for integers $m,n\geq 1$, with basis $\lambda_{1}=(m,0)$ and $\lambda_{2}=(0,n)$ corresponding to $S=C_{m}\times C_{n}$. As a $W$-module it is $\mathbb{Z}\oplus\tilde{\mathbb{Z}}$, and it has Weyl group $T\times C_{2}$
Type 2: $\Lambda_{2}(2m,2n)$ with basis $\lambda_{1}=(m,n)$, $\lambda_{2}=(m,-n)$ for integers $m,n\geq 1$. We have index 2 inclusions
$\Lambda_{1}(2m,2n)\subset\Lambda_{2}(2m,2n)\subset\Lambda_{1}(m,n),$
and the group $S$ has order $2mn$, and $S=C_{2m}\times_{2}C_{2n}$. As a $W$-module $\Lambda_{2}(2m,2n)$ is $\mathbb{Z}W$ and it has Weyl group $T$.
Proof : Every rank 2 lattice $\Lambda$ contains an element on the $x$-axis, and we may choose the smallest positive element $(m,0)$, and similarly $(0,n)$ is the smallest positive point on the $y$-axis. If $\Lambda$ is generated by $(m,0)$ and $(0,n)$ we have a Type 1 lattice. Otherwise there is another element $(a,b)$ with $0<a<m,0<b<n$. Since g(a,b)=(a,b)g(a,b)=(a,-b) we find $(2a,0)$ is in the lattice so $2a=m$. If $2b$ is not a multiple of $n$ this contradicts the choice of $n$. Hence $(a,b)=(m,n)/2$ and we have a Type 2 lattice. ∎
8.C. Sections and conjugacy
Now we need understand the sequence
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
where $i^{S}_{*}$ is the focus of attention (we refer to Section 6 for a summary, including the cohomology groups).
If $S$ is of rank 1, it has identity component either $Z$ or $\tilde{T}$.
In the first case (corresponding to $\Lambda_{-}(n)$), $T^{2}/S$ is a circle with nontrival action so the cohomology of $T^{2}/S$ is in even degrees, and the lift of $\epsilon(G)$ is unique is unique up to conjugacy if it exists. In the split case $\epsilon(G)=0$ so the lift exists. In the non-split case, it lifts if and only if $n$ is even.
In the second case (corresponding to $\Lambda_{+}(m)$), $T^{2}/S$ is a circle with trivial action so the cohomology of $T^{2}/S$ is in odd degrees, so the lift of $\epsilon(G)$ always exists. If $m$ is odd there is a single isomorphism class in the lift, falling into two $G$-conjugacy classes. If $m$ is even there are two isomorphism classes of lifts, each a single $G$-conjugacy class.
Now let us turn to the case when $S$ is finite, so $\Lambda^{S}$ has rank 2.
If $\Lambda^{S}=\Lambda_{1}(m,n)$ is of Type 1, both $\Lambda_{0}$ and $\Lambda_{S}$ are isomorphic to $\mathbb{Z}\oplus\tilde{\mathbb{Z}}$ so that $T^{2}$ and $T^{2}/S$ both have cohomology $\mathbb{Z}/2$ in each positive degree. The map $a^{1}$ is multiplication by $m$ and the map $a^{2}$ is multiplication by $n$. There are therefore four cases.
If $m$ and $n$ are both even, then irrespective of $\epsilon(G)$, there are two lifts to $S$ and each of them forms a single conjugacy class. If $\epsilon(G)=0$ one of the lifts is split and the other is not.
If $m$ and $n$ are both odd, only the split case $\epsilon(G)=0$ lifts. It lifts to a split extension, which falls into two conjugacy classes.
If $m$ is even (so $a^{1}=0$) and $n$ is odd (so $a^{2}$ is an isomorphism) then only the split case $\epsilon(G)=0$ lifts, there is a split lift and a non-split lift, and each forms a single conjugacy class.
If $m$ is odd (so $a^{1}$ is an isomorphism) and $n$ is even (so $a^{2}=0$), then $\epsilon(G)$ always lifts and there is a single extension class of lifts, which falls into two conjugacy classes.
If $\Lambda^{S}=\Lambda_{2}(2m,2n)$ is of Type 2, then as $W$-modules $\Lambda^{0}\cong\mathbb{Z}\oplus\tilde{\mathbb{Z}}$ and $\Lambda^{S}\cong\mathbb{Z}[W]$. Thus $H^{*}(W;T^{2})$ is $\mathbb{Z}/2$ in each positive degree, whilst $H^{*}(W;T^{2}/S)=0$ in positive degrees. The class $\epsilon(G)$ (whichever it is) lifts to $S$ and all lifts are $G$-conjugate.
8.D. Summary
In the split case we have the following picture.
Along the top row we have the rank 1 subgroups $H$ containing $Z$. These are uniquely determined by $S$ and are all split extensions; their lattices are generated by $(0,1),(0,2),(0,3),\ldots$ in turn. The $n$th term is $T\times D_{2n}$.
Along the right hand vertical we have the rank 1 subgroups $H$ containing $\tilde{T}$. For each of the lattices $(1,0),(2,0),(3,0),\cdots$ there are two conjugacy classes of subgroups. If the generator is odd (i.e., $(2n+1,0)$) the lifts are necessarily split but fall into two conjugacy classes. If the generator is even (i.e. $(2n,0)$) one is the split extension and one is the nonsplit extension. The $n$th split term is $C_{n}\times O(2)$, and the non-split term is determined by the fact that the square of the lift of $w$ is the central element of order 2, which may be written $C_{2n}\times_{2}Pin(2)$.
For $1\leq m,n<\infty$, the Type 1 lattice $\Lambda_{1}(m,n)$ corresponds to the subgroup $S$ of order $mn$, giving subgroups $H$ of order $2mn$, two conjugacy classes in each case. If $m$ is even there is a split extension $H_{1}^{s}(m,n)$ (one conjugacy class) and a non-split extension $H_{1}^{ns}(m,n)$ (one conjugacy class). If $m$ is odd there is a single extension class falling into two conjugacy classes $H_{1}^{s,\pm}(m,n)$.
The inclusions of lattices
$\langle(0,n)\rangle\subseteq\Lambda_{1}(m,n)\supseteq\langle(m,0))\rangle$
are both cofree, and the first is cofree and $W$-trivial. They therefore appear as vertical and horizontal inclusions respectively in the diagram.
The lattice $\Lambda_{2}(2m,2n)$ corresponds to the subgroup $S$ of order $2mn$, giving subgroups $H$ of order $4mn$. There is a single conjugacy class and the subgroup $H$ is a split extension $H_{2}(m,n)=S\rtimes W$. We have inclusions of lattices
$\langle(0,2n))\rangle\subseteq\Lambda_{2}(2m,2n)\supseteq\langle(2m,0)\rangle$
(this is the reason for indexing the lattice $\Lambda_{2}(2m,2n)$ using $(2m,2n)$, rather than $(m,n)$). Recalling $\Lambda_{2}(2m,2n)$ is generated by (m,n),(m,n)(m,n),(m,-n) one may check that the inclusions are both cofree and that the first is cofree and $W$-cotrivial. They are therefore represented vertically and horizontally in the picture.
We note that for each finite subgroup $F$ there is a subgroup $\varpi(F)$ with finite Weyl group in which it is cotoral, and $\varpi(F)$ is unique up to conjugacy. We may therefore arrange finite subgroups in cotoral strings leading up to one of the one dimensional subgroups containing $Z$.
${H(\infty,1)}$${H(\infty,2)}$${H(\infty,3)}\textstyle{\cdots}$${H(\infty,\infty)}\textstyle{\vdots}\textstyle{\vdots}\textstyle{\vdots}\textstyle{\vdots}$${H_{1}^{s/ns}(4,1)}$${H_{1}^{s/ns}(4,2),H_{2}(4,2)}$${H_{1}^{s/ns}(4,3)}\textstyle{\cdots}$${H^{s/ns}(4,\infty)}$${H_{1}^{s,\pm}(3,1)}$${H_{1}^{s,\pm}(3,2)}$${H_{1}^{s,\pm}(3,3)}\textstyle{\cdots}$${H^{s,\pm}(3,\infty)}$${H_{1}^{s/ns}(2,1)}$${H_{1}^{s/ns}(2,2),H_{2}(2,2)}$${H_{1}^{s/ns}(2,3)}\textstyle{\cdots}$${H^{s/ns}(2,\infty)}$${H_{1}^{s,\pm}(1,1)}$${H_{1}^{s,\pm}(1,2)}$${H_{1}^{s,\pm}(1,3)}\textstyle{\cdots}$${H^{s,\pm}(1,\infty)}$
Filling in the outer values is rather straightforward.
$H(\infty,n)=T\times D_{2n},H^{s}(m,\infty)=C_{m}\times O(2),H^{ns}(m,\infty)=C_{2m}\times_{2}Pin(2).$
The values for finite subgroups involve more thought, where we have
$H^{s}_{1}(m,n)=C_{m}\times D_{2n},H^{ns}_{1}(m,n)=C_{m}\times_{2}Q_{4n},H_{1}^{s,\pm}(m,n)=C_{m}\times D_{2n},H_{2}(2m,2n)=C_{2m}\times_{2}D_{2n}$
giving the picture.
${T\times D_{2}}$${T\times D_{4}}$${T\times D_{6}}\textstyle{\cdots}$${T\times O(2)}$${(C_{4}\times D_{2}),(C_{4}\times_{2}Q_{4})}$${(C_{4}\times D_{4}),(C_{4}\times_{2}Q_{8}),(C_{4}\times_{2}D_{4})}$${(C_{4}\times D_{6}),(C_{4}\times_{2}Q_{12})}\textstyle{\cdots}$${(C_{4}\times O(2)),(C_{4}\times_{2}Pin(2))}$${(C_{3}\times D_{2})^{2}}$${(C_{3}\times D_{4})^{2}}$${(C_{3}\times D_{6})^{2}}\textstyle{\cdots}$${(C_{3}\times O(2))^{2}}$${(C_{2}\times D_{2}),(C_{2}\times_{2}Q_{4})}$${(C_{2}\times D_{4}),(C_{2}\times_{2}Q_{8}),(C_{4}\times_{2}D_{4})}$${(C_{2}\times D_{6}),(C_{2}\times_{2}Q_{12})}\textstyle{\cdots}$${(C_{2}\times O(2)),(C_{2}\times_{2}Pin(2))}$${(C_{1}\times D_{2})^{2}}$${(C_{1}\times D_{4})^{2}}$${(C_{1}\times D_{6})^{2}}\textstyle{\cdots}$${(C_{1}\times O(2))^{2}}$
In effect we have a diagram of the form
[…]
In the non-split case the diagram is similar, but each of the odd columns is deleted.
In effect the regularity of the pattern is dictated by the pattern of functors $H^{i}(W;\mathbb{T}/(\cdot))$.
8.E. Component structure
We read the component structure off Lemma 8.1 and Lemma 8.2. There is a trivial group over all subgroups on hte right hand vertical (those containing $\tilde{T}$) and all those corresponding to Type 2 lattices. Over all others it is a group of order 2.
9. The case $W=C_{2}$, $\Lambda_{0}=\tilde{\mathbb{Z}}\oplus\tilde{\mathbb{Z}}$
We describe here the space of full subgroups when the $W$-fixed submodule is trivial. The split extension $T^{2}\rtimes W$ with $W$ acting by taking inverses is one example.
9.A. Preamble
In this case $W$ is of order 2, and the toral lattice is $\Lambda_{0}=\tilde{\mathbb{Z}}\oplus\tilde{\mathbb{Z}}$. Thus we have $H^{2}(W;T^{2})=H^{3}(W;\Lambda_{0})=\mathbb{Z}/2\oplus\mathbb{Z}/2$. The split example follows from the action, the three non-split examples are isomorphic but distinguished by which of the three elements of order 2 in $T^{2}$ is the square of an element over the generator of $W$.
9.B. Lattices and Weyl groups
For subgroups we must classify $W$-invariant lattices in $\tilde{\mathbb{Z}}\oplus\tilde{\mathbb{Z}}$ and identify their Weyl groups. We have described the method in Subsection 4.B above. For this lattice the matrix $M$ is given by
$M=\left(\begin{array}[]{cc}-2&0\\ 0&-2\end{array}\right)$
Lemma 9.1.
(i) The unique rank 0 sublattice of $\Lambda_{0}$ is 0, and it corresponds to $S=T^{2}$. It has trivial Weyl group.
(ii) Every rank 1 lattice is $W$-invariant, and $\Lambda_{1}(m,n)$ has basis $\lambda=(m,n)$ where we may suppose $(m,n)$ is a non-zero integer vector with $m\geq 0$. As a $W$-module, the lattice is $\tilde{\mathbb{Z}}$. The lattice $\Lambda^{+}$ is $2\Lambda$ and the Weyl group is of order 2.
Proof : Since $M$ is a scalar matrix, the statements about $\Lambda^{+}$ and the Weyl group are clear. ∎
Lemma 9.2.
All rank 2 lattices are $W$-invariant in $\tilde{\mathbb{Z}}\oplus\tilde{\mathbb{Z}}$. In every case $\Lambda\cong\tilde{\mathbb{Z}}\oplus\tilde{\mathbb{Z}}$, $\Lambda^{+}=2\Lambda$ and the Weyl group is $C_{2}\times C_{2}$.
Proof : Since $M$ is a scalar matrix, the statements about $\Lambda^{+}$ and the Weyl group are clear. ∎
9.C. Sections and conjugacy
Now we need to understand the sequence
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
where $i^{S}_{*}$ is the focus of attention (we refer to Section 6 for a summary, including the cohomology groups).
This time $H^{1}(W;T^{2})=H^{1}(W;T^{2}/S)=0$.
In the non-split case, $\epsilon(G)$ lifts to $S$ if and only if $S$ contains the element of $T^{2}$ which is the square of an element above the generator of $W$, and if it exists, it is unique up to $G$-conjugacy.
9.D. Summary
We consider a 2-dimensional toral group $G$ with two components and $\Lambda_{0}=\tilde{\mathbb{Z}}\oplus\tilde{\mathbb{Z}}$. Up to isomorphism, there are only two of these: one is the split extension and the other is a non-split extension (there are three non-split extensions, but they are abstractly isomorphic).
9.D.1. The space of full subgroups of the split extension
If $G$ is a split extension then any full subgroup is also a split extension, and the subgroups are in bijective correspondence to subgroups $S$ of $T^{2}$. As usual we will consider the corresponding subgroups $\Lambda^{S}$ of $\Lambda^{0}$. Because $W$ acts as negation, there is no restriction on $\Lambda^{S}$.
Now consider the space $\mathrm{Sub}(T^{2})$. There is one 2-dimensional group. Then the 1-dimensional groups $S$ correspond to 1-dimensional lattices $\Lambda^{S}$. We may think of the 1-dimensional lattices as lying over $\mathbb{P}^{1}(\mathbb{Q})$ with $S$ mapping to $\Lambda^{S}\otimes\mathbb{Q}$. The fibres of this map can be identified with the positive integers, as discussed at the end of Remark 5.7. The 0-dimensional groups $F$ correspond to 2-dimensional lattices $\Lambda^{F}$.
9.D.2. The space of full subgroups of the non-split extension
If $G$ is a non-split extension then one of the three elements of $T^{2}$ of order 2 is distinguished, because it is a square of an element outside $T^{2}$. Writing $x_{e}$ for the distinguished element, the full subgroups are in bijective correspondence to subgroups of $T^{2}$ containing $x_{0}$. The corresponding subgroups of $\Lambda^{S}$ are precisely those lying in the index 2 sublattice $\Lambda^{e}\subseteq\Lambda^{0}$ corresponding to $x_{e}$. Because $W$ acts as negation, there is no further restriction on $\Lambda^{S}$.
The space of subgroups can be described exactly as in the split case but with $\Lambda^{0}$ replaced by $\Lambda^{e}$.
9.D.3. Component structure
The Weyl group of a subgroup $S$ is the dual of $\Lambda^{S}/2\Lambda^{S}$, which is $C_{2}\times C_{2}$ for finite subgroups, $C_{2}$ for 1-dimensional subgroups and is trivial for $G$. This also gives the component structure since if $F\subset S$ we find $\Lambda^{F}\supset\Lambda^{S}$, and there is a map $(\Lambda^{F}/2\Lambda^{F})\longleftarrow(\Lambda^{S}/2\Lambda^{S})$ whose dual is the map in the component structure.
10. The case $W=C_{2}$, $\Lambda_{0}=\mathbb{Z}W$
We describe here the space of full subgroups when the toral lattice is the group ring; this is realised as the normalizer of the maximal torus in $U(2)$.
10.A. Preamble
We have $W$ of order 2 and the toral lattice is $\Lambda_{0}=\mathbb{Z}W$. In this case $H^{2}(W;T^{2})=H^{3}(W;\Lambda_{0})=0$, and all such extensions split. To be concrete $W$ acts on $T^{2}$ by exchanging factors: this example is realised by the normalizer of the maximal torus in $U(2)$. The diagonal circle $Z$ is central and the antidiagonal $\tilde{T}$ consists of pairs $(z,z^{-1})$.
10.B. Lattices and Weyl groups
For subgroups we must classify $W$-invariant lattices in $\mathbb{Z}[W]$ and identify their Weyl groups. We have described the method in Subsection 4.B above. We choose the basis of $\mathbb{Z}W$ given by $1,w$. For this lattice the matrix $M$ is given by
$M=\left(\begin{array}[]{cc}-1&1\\ 1&-1\end{array}\right)$
Lemma 10.1.
(i) The unique rank 0 sublattice of $\Lambda^{0}$ is 0, and it corresponds to $S=T^{2}$. It has trivial Weyl group.
(ii) The rank 1 lattices in $\Lambda^{0}=\mathbb{Z}W$ are of two types
(a) The lattice $\Lambda_{+}(m)$ with basis $\lambda=(m,m)=m(1+w)$ for an integer $m\geq 1$. As a $W$-module, this is $\mathbb{Z}$. It corresponds to $S$ with $S_{e}=\tilde{T}$ and $S\cong C_{m}\times\tilde{T}$ with $\tilde{T}$. The Weyl group is and has Weyl group $Z/C_{m}$.
(b) The lattice $\Lambda_{-}(m)$ with basis $\lambda=(m,-m)=m(1-w)$ for an integer $y\geq 1$. As a $W$-module, this is $\tilde{\mathbb{Z}}$. This corresponds to $S$ with $S_{e}=Z$ and $S\cong Z\times C_{m}$, and has Weyl group of order 2.
Proof : If $\Lambda$ is a rank 1 lattice generated by $\lambda$ we need $g\lambda=\pm\lambda$. ∎
Lemma 10.2.
There are two families of rank 2 $W$-invariant lattices in $\mathbb{Z}W$.
Type 1: $\Lambda_{1}(m,n)$ with basis $\lambda_{1}=(m,m)=m(1+w)$, $\lambda_{2}=(n,-n)=n(1-w)$ for integers $m\geq 1,n\geq 1$, which is $\mathbb{Z}\oplus\tilde{\mathbb{Z}}$ as a $W$-module. We have containments
$\Lambda_{1}(m,n)\subset\Lambda_{1}(1,1)\subset\Lambda_{0}=\mathbb{Z}W.$
The corresponding subgroup $S$ has order $2mn$ and is $C_{2m}\times_{2}C_{2n}$. The Weyl group of $H$ is $T\times C_{2}$.
Type 2: If $m+n$ and mnm-n are even we have $\Lambda_{2}(m,n)$ which contains $\Lambda_{1}(m,n)$ as a sublattice of index 2. It has basis $\lambda_{1}=((m+n)/2,(m-n)/2)=((m+n)+(m-n)w)/2$ and $\lambda_{2}=w\lambda_{1}=((m-n)/2,(m+n)/2)=((m-n)+(m+n)w)/2$ and is $\mathbb{Z}W$ as a $W$-module. We have containments
$\Lambda_{1}(m,n)\subset\Lambda_{2}(m,n).$
The corresponding subgroup $S$ has order $mn$ and contains cyclic groups of order $m$ and $n$ (the presentation is $\left(\begin{array}[]{cc}m&\frac{m-n}{2}\\ 0&n\end{array}\right)$). The subgroup $H$ has Weyl group $T$.
Proof : The eigenspaces of $\mathbb{Z}W$ are generated by $(1+w)$ and (1w)(1-w). If a lattice $\Lambda$ is generated by its intersections with $\mathbb{Z}W_{+}$ and $\mathbb{Z}W_{-}$, then it is of Type 1. Otherwise there will be an additional element, and we argue as in Lemma 8.2 to see that the lattice must be of Type 2. ∎
10.C. Sections and conjugacy
Now we need understand the sequence
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
where $i^{S}_{*}$ is the focus of attention. Since $H^{i}(W;\mathbb{T})=0$ for $i\neq 0$, we find $H^{1}(W;T^{2}/S)\cong H^{2}(W;S)$ (we refer to Section 6 for a summary, including the cohomology groups).
If $S$ is of rank 1, it has identity component either $Z$ or $\tilde{T}$.
In the first case (corresponding to $\Lambda_{-}(n)$), $T^{2}/S$ is a circle with nontrival action, so the cohomology of $T^{2}/S$ is in even degrees, and $H(S,\sigma)$ must be the split extension, and this is unique up to $G$-conjugacy.
In the second case (corresponding to $\Lambda_{+}(m)$), $T^{2}/S$ is a circle with trivial action so the cohomology of $T^{2}/S$ is in odd degrees so $H^{2}(W;S)=\mathbb{Z}/2$ and there is a split and a non-split extension, each unique up to $G$-conjugacy.
Turning to the case when $S$ is finite, there are two cases. If the lattice is Type 1, $H^{i}(W;T^{2}/S)=\mathbb{Z}/2$ for each $i>0$. There is a split lift and a non-split lift, each unique up to conjugacy. If the lattice is of Type 2, then $H^{i}(W;T^{2}/S)=0$ for each $i>0$. The split extension is the only lift and it forms a single $G$-conjugacy class.
10.D. Summary
In general terms this is exactly like the case $\mathbb{Z}\oplus\tilde{\mathbb{Z}}$, but the lattices are more subtle. The general picture in terms of lattices is not hard to work out, but the structure of the individual groups needs some care.
Along the top row we have the rank 1 subgroups $H$ containing the centre $Z$ (i.e., elements $(z,z)\in T^{2}$). These are uniquely determined by $S$ and are all split extensions; their lattices are generated by $(1-w),2(1-w),3(1-w),\ldots$ in turn. The $n$th term is $T\times D_{2n}$.
Along the right hand vertical we have the rank 1 subgroups $H$ containing the anticentre $\tilde{T}$ (i.e., elements $(z,z^{-1})\in T^{2}$). For each of the lattices $(1+w),2(1+w),3(1+w),\cdots$ there are two conjugacy classes, one is the split extension and one is the nonsplit extension. The $n$th split term is $C_{n}\times O(2)$, and the non-split term is determined by the fact that the square of the lift of $w$ is the central element of order 2, which may be written $C_{2n}\times_{2}Pin(2)$.
For $1\leq m,n<\infty$, the Type 1 lattice $\Lambda_{1}(m,n)$ corresponds to the subgroup $S$ of order $2mn$, giving subgroups $H$ of order $4mn$. There is a split extension $H_{1}^{s}(m,n)$ and a non-split extension $H_{1}^{ns}(m,n)$. The inclusions of lattices
$\langle n(1-n)\rangle\subseteq\Lambda_{1}(m,n)\supseteq\langle m(1+w)\rangle$
are both cofree, and the first is cofree and $W$-cotrivial. They therefore appear as vertical and horizontal inclusions respectively in the diagram.
The lattice $\Lambda_{2}(m,n)$ corresponds to the subgroup $S$ of order $mn$, giving subgroups $H$ of order $2mn$. There is a single conjugacy class and the subgroup $H$ is a split extension $H_{2}(m,n)=S\rtimes W$. We have inclusions of lattices
$\langle n(1-w)\rangle\subseteq\Lambda_{2}(m,n)\supseteq\langle m(1+w)\rangle$
and one may check that the inclusions are both cofree and that the first is cofree and $W$-cotrivial. They are therefore represented vertically and horizontally in the picture.
We note that for each finite subgroup $F$ there is a subgroup $\varpi(F)$ with finite Weyl group in which it is cotoral, and $\varpi(F)$ is unique up to conjugacy. We may therefore arrange finite subgroups in cotoral strings leading up to one of the one dimensional subgroups containing $Z$.
${H(\infty,1)}$${H(\infty,2)}$${H(\infty,3)}\textstyle{\cdots}$${H(\infty,\infty)}\textstyle{\vdots}\textstyle{\vdots}\textstyle{\vdots}\textstyle{\vdots}$${H_{1}^{s/ns}(4,1)}$${H_{1}^{s/ns}(4,2),H_{2}(4,2)}$${H_{1}^{s/ns}(4,3)}\textstyle{\cdots}$${H^{s/ns}(4,\infty)}$${H_{1}^{s/ns}(3,1),H_{2}(3,1)}$${H_{1}^{s/ns}(3,2)}$${H_{1}^{s/ns}(3,3),H_{2}(3,3)}\textstyle{\cdots}$${H^{s/ns}(3,\infty)}$${H_{1}^{s/ns}(2,1)}$${H_{1}^{s/ns}(2,2),H_{2}(2,2)}$${H_{1}^{s/ns}(2,3)}\textstyle{\cdots}$${H^{s/ns}(2,\infty)}$${H_{1}^{s/ns}(1,1),H_{2}(1,1)}$${H_{1}^{s/ns}(1,2)}$${H_{1}^{s/ns}(1,3),H_{2}(1,3)}\textstyle{\cdots}$${H^{s/ns}(1,\infty)}$
Filling in the outer values is rather straightforward.
$H(\infty,n)=T\times D_{2n},H^{s}(m,\infty)=C_{m}\times O(2),H^{ns}(m,\infty)=C_{2m}\times_{2}Pin(2).$
The values for finite subgroups involve more thought, where we have
$H^{s}_{1}(m,n)=C_{2m}\times_{2}D_{4n},H^{ns}_{1}(m,n)=C_{2m}\times_{2}Q_{4n},H_{2}(m,n)=(\Lambda_{0}/\Lambda_{2}(m,n))\rtimes W$
giving the picture.
${T\times D_{2}}$${T\times D_{4}}$${T\times D_{6}}\textstyle{\cdots}$${T^{2}\rtimes W}\textstyle{\vdots}\textstyle{\vdots}\textstyle{\vdots}\textstyle{\vdots}$${(C_{8}\times_{2}D_{4}),(C_{8}\times_{2}Q_{4})}$${(C_{8}\times_{2}D_{8}),(C_{8}\times_{2}Q_{8})}$${(C_{8}\times_{2}D_{12}),(C_{8}\times_{2}Q_{12})}\textstyle{\cdots}$${(C_{4}\times O(2)),(C_{8}\times_{2}Pin(2))}$${(C_{6}\times_{2}D_{4}),(C_{6}\times_{2}Q_{4}),H_{2}(3,1)}$${(C_{6}\times_{2}D_{8}),(C_{6}\times_{2}Q_{8})}$${(C_{6}\times_{2}D_{12}),(C_{6}\times_{2}Q_{12}),H_{2}(3,3)}\textstyle{\cdots}$${(C_{3}\times O(2)),(C_{6}\times_{2}Pin(2))}$${(C_{4}\times_{2}D_{4}),(C_{4}\times_{2}Q_{4})}$${(C_{4}\times_{2}D_{8}),(C_{4}\times_{2}Q_{8}),H_{2}(2,2)}$${(C_{4}\times_{2}D_{12}),(C_{4}\times_{2}Q_{12})}\textstyle{\cdots}$${(C_{2}\times O(2)),(C_{4}\times_{2}Pin(2))}$${(C_{2}\times_{2}D_{4}),(C_{2}\times_{2}Q_{4}),H_{2}(1,1)}$${(C_{2}\times_{2}D_{8}),(C_{2}\times_{2}Q_{8})}$${(C_{2}\times D_{12}),(C_{2}\times Q_{12}),H_{2}(1,3)}\textstyle{\cdots}$${(C_{1}\times O(2)),(C_{2}\times_{2}Pin(2))}$
In effect we have a diagram of the form
${\mathbb{N}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{N}^{0}}$${2\mathbb{N}^{2}+\Lambda\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${2\mathbb{N}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
In the non-split case the diagram is similar, but each of the odd columns is deleted.
In effect the regularity of the pattern is dictated by the pattern of functors $H^{i}(W;\mathbb{T}/(\cdot))$.
10.E. Component structure
We read the component structure off Lemma 10.1 and Lemma 10.2. There is a trivial group over all subgroups on hte right hand vertical (those containing $\tilde{T}$) and all those corresponding to Type 2 lattices. Over all others it is a group of order 2.
10.F. The view inside $U(2)$ torus normalizer
To start with, we choose the maximal torus $T^{2}$ to consist of the diagonal matrices. The Weyl group $W$ of order 2 acts on this by exchanging the entries. We write $Z$ for scalar matrices, and note that it is the centre, and it is the factor $T$ in the decomposition $U(2)=T\times_{C_{2}}SU(2)$. We write $\tilde{T}$ for $T^{2}\cap SU(2)$ (the matrices $\mathrm{diag}(\lambda,\lambda^{-1})$). This is a maximal torus for $SU(2)$, and $L\tilde{T}$ is the 1-1 eigenspace for the action of $W$. We note that $T^{2}$ is generated by $Z$ and $\tilde{T}$ but $Z\cap\tilde{T}$ is of order 2, so that $T^{2}=Z\times_{C_{2}}\tilde{T}$.
We will also write $B\overline{\times}C$ for the image of $B\times C$ in $U(2)$ if $B$ is a subgroup of $Pin(2)$ or $O(2)$ and $C$ is a subgroup of $T$, noting that when $B$ is cyclic of order $\geq 2$ the two possible interpretations are different but conjugate. If $B$ and $C$ contain $C_{2}$ we have $B\overline{\times}C=B\times_{C_{2}}C$, but otherwise the projection $SU(2)\times T\longrightarrow U(2)$ will give an isomorphism $B\times C\cong B\overline{\times}C$; the $\overline{\times}$ notation allows us to treat these cases together.
Considering the maximal torus normalizer in $SU(2)$ we see $G$ is generated by $T^{2}$ and the element $\tilde{w}:=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)$ of order $4$. This shows $G$ contains $Pin(2)$, and in fact $G=Pin(2)\times_{C_{2}}T$.
On the other hand, we note that the subgroup $G$ is also generated by the $T^{2}$ together with the element $w:=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)$ of order $2$. As such it is a split extension $T^{2}\rtimes C_{2}$. Indeed, this makes it apparent that it contains $O(2)$ (which was obvious in any case by looking at real matrices in $U(2)$). It is visible that $O(2)$ contains the central subgroup $C_{2}=\{\pm I\}$ and so we have
$T\times_{C_{2}}O(2)=N_{U(2)}(T^{2})=T\times_{C_{2}}Pin(2).$
11. The case $W=C_{2}\times C_{2}$ acting on $\Lambda_{0}$ as an $A_{1}\times A_{1}$-lattice
11.A. Preamble
This example is the normalizer of the maximal torus inside a group of local type $A_{1}\times A_{1}$ (such as $SU(2)\times SU(2)$). In other words $\Lambda_{0}$ is the $A_{1}\times A_{1}$-lattice and taken with the Lie theoretic Weyl group $W=D_{4}\cong C_{2}\times C_{2}$ acting on the lattice. In concrete terms $D_{4}$ acts on $\Lambda_{0}=\tilde{\mathbb{Z}}_{x}\oplus\tilde{\mathbb{Z}}_{y}$ by reflection in the $x$-axis and the $y$-axis (where $\tilde{\mathbb{Z}}_{x},\tilde{\mathbb{Z}}_{y}$ consist of integral points on the $x$ and $y$ axes). We have $H^{3}(W;\Lambda_{0})\cong(\mathbb{Z}/2)^{4}$, where the large number of cases arises because of the choice of which element of order 2 in $W$ lifts to an element of order 4, and which element of order 2 in $T^{2}$ is a square in $G$.
11.B. Lattices and Weyl groups
For subgroups we must classify $W$-invariant lattices in $\Lambda^{0}$ and identify their Weyl groups. We have described the method in Subsection 4.B above. For this lattice the group $W$ has two generators, and the two matrices $M$ are given by
$M_{1}=\left(\begin{array}[]{cc}0&0\\ 0&-2\end{array}\right),M_{2}=\left(\begin{array}[]{cc}-2&0\\ 0&0\end{array}\right)$
for reflection in the $x$ axis.
Lemma 11.1.
. (i) The unique rank 0 sublattice of $\Lambda^{0}$ is 0, and it corresponds to $S=T^{2}$. It has trivial Weyl group.
(ii) The rank 1 lattices are $\Lambda_{x}(m)=\langle(m,0)\rangle$ or $\Lambda_{y}(m)=\langle(0,m)\rangle$. These have Weyl groups $C_{2}$.
Lemma 11.2.
If we pick an element of order 2 in $W$, we have an instance of the case $\Lambda_{0}=\mathbb{Z}\oplus\tilde{\mathbb{Z}}$ in Section 8. Every invariant lattice from that case is also invariant under $W=D_{4}$, so they are
$\Lambda_{1}(m,n)=\langle(m,0),(0,n)\rangle\mbox{ or }\Lambda_{2}(m,n)=\langle(m,n),(m,-n)\rangle.$
They all have Weyl group $C_{2}\times C_{2}$. We have containments
$\Lambda_{1}(2m,2n)\subseteq\Lambda_{2}(m,n)\subseteq\Lambda_{1}(m,n).$
11.C. Sections and conjugacy
Now we need understand the sequence
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
where $i^{S}_{*}$ is the focus of attention (we refer to Section 6 for a summary, including the cohomology groups).
Each of the two rank 1 lattices is an inflation of $\tilde{\mathbb{Z}}$ from a quotient group of order 2. Hence $H^{1}(W;T^{2}/S)\cong H^{2}(W;\tilde{\mathbb{Z}})=\mathbb{Z}/2$ and $H^{2}(W;T^{2}/S)\cong H^{3}(W;\tilde{\mathbb{Z}})=(\mathbb{Z}/2)^{2}$. This can be calculated directly, or the answer can be obtained by considering both Serre spectral sequences for the projection $D_{4}\longrightarrow C_{2}$ (in low degrees, comparing these gives the differential).
In each case half of the classes in $H^{2}(W;T^{2})$ lift automatically, and the other half lift if $m$ is even. Each lift is unique and the extension splits into two $G$-conjugacy classes.
Now consider the case when $S$ is finite, and $\Lambda$ has rank 2. The Type 1 lattices are isomorphic to $\Lambda^{0}$ and hence of the form $\tilde{\mathbb{Z}}_{x}\oplus\tilde{\mathbb{Z}}_{y}$, with cohomology as above. The behaviour depends on the parities of $m$ and $n$. If both are odd, only the split extension lifts; it lifts uniquely and then splits into four conjugacy classes. If both $m$ and $n$ are even, all classes lift each one in four ways, and each forms a single conjugacy class. If one is even and one is odd, half the classes lift; each lifts in two ways and each of those splits into two conjugacy classes.
The Type 2 lattices are isomorphic to $\mathrm{ind}_{C_{2}}^{W}\tilde{\mathbb{Z}}$, and hence have $H^{1}(W;T^{2}/S)\cong H^{2}(W;\Lambda^{S})=H^{2}(C_{2};\tilde{\mathbb{Z}})=0$ and $H^{2}(W;T^{2}/S)\cong H^{3}(W;\Lambda^{S})=H^{3}(C_{2};\tilde{\mathbb{Z}})=\mathbb{Z}/2$. If $m+n$ and mnm-n are both even, all classes lift, and otherwise exactly half othem do. Those classes that lift, do so uniquely, to a single conjugacy class.
11.D. Summary
Once again, the pattern can be layed out with subgroups $H^{\lambda}(m,n)$ with $1\leq m,n\leq\infty$.
In the split case there are 2 subgroups $H^{s/ns}(\infty,n)$ or $H^{s/ns}(1,\infty)$. When $m,n$ are finite there are 4 subgroups for each pair $(m,n)$ from Type 1 lattices and an additional one from Type 2.
In each nonsplit case there are 2 subgroups $H(\infty,n)$ when $n$ is even and 0 or 2 subgroups otherwise, and similarly there are 2 subgroups $H(m,\infty)$ when $n$ is even and 0 or $2$ subgroups otherwise depending on $m$. When $m,n$ are finite there are 4 subgroups for certain pairs $(m,n)$ including those with $m$ and $n$ even from Type 1 lattices and an additional 1 from Type 2 from various pairs $m$, $n$ including those where $m+n$ and mnm-n are even.
12. The case $W=C_{2}\times C_{2}$ acting on $\Lambda_{0}=\mathbb{Z}\oplus\mathbb{Z}$ via diagonal reflections
12.A. Preamble
We take $\Lambda_{0}$ to be $\mathbb{Z}\oplus\mathbb{Z}$, and we take $W=C_{2}\times C_{2}$ and let it act on $\mathbb{Z}\oplus\mathbb{Z}$ by reflection in the lines $x=\pm y$. This is one of the two subgroups of $D_{8}$ acting on the $B_{2}$ lattice, the other having been covered as $A_{1}\times A_{1}$ in Section 11. In this case, restricting to either factor $C$ gives $\mathbb{Z}C$, and $H^{3}(W;\Lambda_{0})\cong\mathbb{Z}/2,$ so there is one split and one non-split example.
12.B. Lattices and Weyl groups
For subgroups we must classify $W$-invariant lattices in $\Lambda^{0}$ and identify their Weyl groups. We have described the method in Subsection 4.B above. For this lattice the group $W$ has two generators, and the two matrices $M$ are given by
$M_{1}=\left(\begin{array}[]{cc}-1&1\\ 1&-1\end{array}\right),M_{2}=\left(\begin{array}[]{cc}-1&-1\\ -1&-1\end{array}\right)$
Lemma 12.1.
. (i) The unique rank 0 sublattice of $\Lambda^{0}$ is 0, and it corresponds to $S=T^{2}$. It has trivial Weyl group.
(ii) The rank 1 lattices are $\Lambda_{+}(m)=\langle(m,m)\rangle$ or $\Lambda_{-}(m)=\langle(m,-m)\rangle$. These have Weyl groups $C_{2}$.
Lemma 12.2.
There are two families of rank 2 $W$-invariant lattices.
Type 1: $\Lambda_{1}(m,n)$ with basis $\lambda_{1}=(m,m)$, $\lambda_{2}=(n,-n)$ for integers $m\geq 1,n\geq 1$, which is a sum $\mathbb{Z}_{+}\oplus\mathbb{Z}_{-}$ as a $W$-module. We have containments
$\Lambda_{1}(m,n)\subset\Lambda_{1}(1,1)\subset\Lambda_{0}.$
The Weyl group of $H$ is $C_{2}\times C_{2}$.
Type 2: If $m+n$ and mnm-n are even we have $\Lambda_{2}(m,n)$ which contains $\Lambda_{1}(m,n)$ as a sublattice of index 2. It has basis $\lambda_{1}=((m+n)/2,(m-n)/2)$ and $\lambda_{2}=((m-n)/2,(m+n)/2)$. We have containments
$\Lambda_{1}(m,n)\subset\Lambda_{2}(m,n).$
The subgroup $H$ has Weyl group $C_{2}\times C_{2}$.
Proof : Any lattice is invariant under a reflection so it is one of the form $\Lambda_{1}(m,n)$ or $\Lambda_{2}(m,n)$ from Section 10, and we see that these are in fact also $W$-invariant. ∎
12.C. Sections and conjugacy
Now we need understand the sequence
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
where $i^{S}_{*}$ is the focus of attention (we refer to Section 6 for a summary, including the cohomology groups).
Consider the rank 1 lattices which are modules $m\tilde{\mathbb{Z}}$ inflated from $W/C$ where $C$ is the subgroup of order 2 corresponding to the line $x=y$. We may see that $H^{2}(W;\tilde{\mathbb{Z}})\neq 0$ by thinking of quotient by a complementary subgroup. Hence we see that in the extension $1\longrightarrow C\longrightarrow W\longrightarrow W/C\longrightarrow 0$ that $d^{3}$ is zero from $E_{2}^{0,2}$ and the map $\Lambda_{0}\longrightarrow\Lambda_{+}(1)^{\vee}$ is nonzero in $H^{3}$, and the general map is by $m$. When the lift exists there are two extensions, each forming a single conjugacy class. A precisely similar analysis applies to $\Lambda_{-}(m)$.
Now consider the case when $S$ is finite, and $\Lambda$ has rank 2. The Type 1 lattices are isomorphic to $\mathbb{Z}_{+}\oplus\mathbb{Z}_{-}$. If $m$ or $n$ is odd then the map is injective and only the split extension lifts. If $m$ and $n$ are even both extensions lift. Whenever $\epsilon(G)$ lifts, there are four extension classes, each a single conjugacy class.
For Type 2 lattices we have the inclusion $\Lambda_{2}(m,n)\longrightarrow\Lambda_{2}(1,1)=\Lambda^{0}$, so the induced map is multiplication by $m$ mod 2 (which is equal to $n$ mod 2). If $m$ is odd, only the split extension lifts. If $m$ is even both extensions lift. In either case, the lifts are unique.
12.D. Summary
Once again, the pattern can be layed out with subgroups $H^{\lambda}(m,n)$ with $1\leq m,n\leq\infty$.
In the split case there are 2 subgroups $H^{s/ns}(\infty,n)$ or $H^{s/ns}(1,\infty)$. When $m,n$ are finite there are 4 subgroups for each pair $(m,n)$ from Type 1 lattices and if $m+n$ and mnm-n are event there is an additional one from Type 2).
In the nonsplit case there are 2 subgroups $H^{s/ns}(\infty,n)$ when $n$ is even, and 2 subgroups $H^{s/ns}(m,\infty)$ when $m$ is even. When $m,n$ are finite there are 4 subgroups for each pair $(m,n)$ with $m$ and $n$ even from Type 1 lattices and an additional 1 from Type 2.
13. The case $W=C_{4}$ acting on $\Lambda_{0}=\mathbb{Z}\oplus\mathbb{Z}$ by quarter turn rotation
13.A. Preamble
We take $W$ to be cyclic of order 4, acting on $\Lambda_{0}=\mathbb{Z}^{2}$ via a quarter turn. As a representation this is $\mathrm{ind}_{C_{2}}^{C_{4}}\tilde{\mathbb{Z}}$, and hence $H^{3}(W;\Lambda_{0})\cong\mathbb{Z}/2$ so there is a split and a non-split group.
13.B. Lattices and Weyl groups
For subgroups we must classify $W$-invariant lattices in $\Lambda^{0}$ and identify their Weyl groups. We have described the method in Subsection 4.B above. For this lattice the matrix $M$ is given by
$M=\left(\begin{array}[]{cc}-1&-1\\ 1&-1\end{array}\right)$
for a quarter turn.
Lemma 13.1.
(i) The unique rank 0 sublattice of $\Lambda^{0}$ is 0, and it corresponds to $S=T^{2}$. It has trivial Weyl group.
(ii) There are no rank 1 lattices.
Lemma 13.2.
The $C_{4}$-invariant rank 2 lattices are
$\Lambda_{1}(m,m)=\langle(m,0),(0,m)\rangle\mbox{ and }\Lambda_{2}(m,m)=\langle(m,m),(m,-m)\rangle.$
They all have Weyl group $C_{2}$. We have containments
$\Lambda_{1}(2m,2m)\subseteq\Lambda_{2}(m,m)\subseteq\Lambda_{1}(m,m).$
13.C. Sections and conjugacy
Now we need understand the sequence
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
where $i^{S}_{*}$ is the focus of attention (we refer to Section 6 for a summary, including the cohomology groups).
There are no rank 1 lattices so we immediately consider the case when $S$ is finite, and $\Lambda$ has rank 2. The Type 1 lattices are isomorphic to $\Lambda^{0}$ and hence of the form $\mathrm{ind}_{C_{2}}^{C_{4}}\tilde{\mathbb{Z}}$, with cohomology as above. The behaviour depends on the parity of $m$. If $m$ odd, only the split extension lifts, but then lifts uniquely. If $m$ is even, all lift and each forms a single conjugacy class.
The Type 2 lattices are also isomorphic to $\mathrm{ind}_{C_{2}}^{C_{4}}\tilde{\mathbb{Z}}$, and hence those classes that lift do so uniquely. The slight surprise is that $a^{2}$ is zero in all cases. To see this we note that the inclusion $\Lambda_{2}(m,m)\subseteq\Lambda_{0}$ factors through $\Lambda_{2}(1,1)\subseteq\Lambda_{1}(1,1)=\Lambda^{0}$. For this we make a calculation. Using the standard periodic resolution, if $g^{2}=-1$ on $\Lambda$, $H^{2i+1}(W;\Lambda)=\Lambda/(1-g)$, and we see the dual of the inclusion $\Lambda_{2}(1,1)\subseteq\Lambda_{1}(1,1)$ induces the zero map.
13.D. Summary
For the split group $G=T^{2}\rtimes C_{4}$ we find $\mathfrak{X}_{G}$ is the one point compactification of
$\mathcal{F}=\{H_{1}(m)\;|\;m\geq 1\}\cup\{H_{2}(m)\;|\;m\geq 1\},$
and the component structure is constant at $C_{2}$ on $\mathcal{F}$. The groups $H_{1}(m)=T^{2}\rtimes C_{4}$ and the groups $H_{2}(m)$ are also split extensions.
For the nonsplit group $G=T^{2}\cdot C_{4}$ we find $\mathfrak{X}_{G}$ is the one point compactification of
$\mathcal{F}=\{H_{1}(2m)\;|\;m\geq 1\}\cup\{H_{2}(m)\;|\;m\geq 1\}$
and the component structure is constant at $C_{2}$ on $\mathcal{F}$. The groups are $H_{1}(m)=T^{2}[m]\cdot C_{4}$, and none of the $H_{i}(m)$ are split extensions.
14. The case $W=D_{8}$ acting on $\Lambda_{0}$ as a $B_{2}$-lattice
14.A. Preamble
This example is the normalizer of the maximal torus inside a group of local type $B_{2}$ (such as $Sp(2)$). In other words $\Lambda_{0}$ is the $B_{2}$-lattice and taken with the Lie theoretic Weyl group $W=D_{8}$ acting on the lattice. In concrete terms $D_{8}$ acts on $\Lambda_{0}=\mathbb{Z}\oplus\mathbb{Z}$ generated by reflection in the axes and diagonals and $H^{2}(W;\Lambda_{0})=\mathbb{Z}/2j,H^{3}(W;\Lambda_{0})=(\mathbb{Z}/2)^{2}$. We will focus the analysis on the split extension. The case of the normalizer of the maximal torus in $Sp(2)$ will be treated elsewhere.
14.B. Lattices and Weyl groups
For subgroups we must classify $W$-invariant lattices in $\Lambda^{0}$ and identify their Weyl groups. We have described the method in Subsection 4.B above. For this lattice the group $W$ has two generators, and the two matrices $M$, for rotation and reflection respectively, are given by
$M_{1}=\left(\begin{array}[]{cc}-1&-1\\ 1&-1\end{array}\right),M_{2}=\left(\begin{array}[]{cc}0&0\\ 0&-2\end{array}\right)$
Lemma 14.1.
. (i) The unique rank 0 sublattice of $\Lambda^{0}$ is 0, and it corresponds to $S=T^{2}$. It has trivial Weyl group.
(ii) There are no rank 1 $W$-invariant sublattices.
Lemma 14.2.
The $W$-invariant sublattices are
$\Lambda_{1}(m,m)=\langle(m,0),(0,m)\rangle\mbox{ and }\Lambda_{2}(m,m)=\langle(m,m),(m,-m)\rangle.$
They all have Weyl group $C_{2}$. We have containments
$\Lambda_{1}(2m,2m)\subseteq\Lambda_{2}(m,m)\subseteq\Lambda_{1}(m,m).$
Proof : Any $W$-invariant lattice is invariant under $C_{4}$ so by Lemma 13.2 the only possibilities are the named lattices, each of them easily seen to be $W$-invariant. Similarly the condition imposed by $M_{1}$ implies that required for $M_{2}$. ∎
14.C. Sections and conjugacy
Now we need understand the sequence
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
where $i^{S}_{*}$ is the focus of attention (we refer to Section 6 for a summary, including the cohomology groups).
Now consider the case when $S$ is finite, and $\Lambda$ has rank 2. We have a commutative square
${H^{i}(W;\Lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{i}(W;\Lambda_{S})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{i}(C_{4};\Lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{i}(C_{4};\Lambda_{S})}$
and for $i=1$ the vertical maps are isomorphisms (it is zero for $i=2,3$). We then claim that the map $H^{i}(W;\Lambda_{0})\longrightarrow H^{i}(W;\Lambda_{S})$ for $i=2,3$ is the same as for $i=1$. We thus refer to the behaviour in Section 13.
For Type 1 lattices $\Lambda_{1}(m,m)$ The behaviour depends on the parity of $m$. If $m$ is odd, only the split extension lifts; there is a unique extension class of lifts, which splits into two conjugacy classes. If $m$ is even, there are again two conjugacy classes of lifts, but now they are not isomorphic as extensions. For Type 2 lattices, there are precisely two conjugacy classes of lifts in each case, and they are not isomorphic as extensions.
14.D. Summary
For the split group $G=T^{2}\rtimes D_{8}$ we find $\mathfrak{X}_{G}$ is the one point compactification of
$\mathcal{F}=\{H_{1}(m),H^{\prime}_{1}(m)\;|\;m\geq 1\}\cup\{H_{2}(m),H^{\prime}_{2}(m)\;|\;m\geq 1\},$
and the component structure is constant at $C_{2}$ on $\mathcal{F}$. Here $H_{1}(m)=T^{2}[m]\rtimes D_{8}$, and we have $H_{1}^{\prime}(m)=T^{2}\cdot D_{8}$ is a split extensions if and only if $m$ is odd. Similarly, $H_{2}(m)$ is a split extension and $H^{\prime}_{2}(m)$ is never a split extension.
For the nonsplit group $G=T^{2}\cdot D_{8}$ we find $\mathfrak{X}_{G}$ is the one point compactification of
$\mathcal{F}=\{H_{1}(2m),H^{\prime}_{1}(2m)\;|\;m\geq 1\}\cup\{H_{2}(m),H^{\prime}_{2}(m)\;|\;m\geq 1\}$
and the component structure is constant at $C_{2}$ on $\mathcal{F}$. In this case none of the the finite subgroups is a split extension and the groups and their primed counterparts are not isomorphic as extensions.
15. The case $W=C_{3}$, $\Lambda_{0}=\mathbb{Z}[\omega]$
15.A. Preamble
Let $W=\langle g\rangle$ be of order 3 and the toral lattice is $\Lambda_{0}=\mathbb{Z}[\omega]$ where $\omega=e^{2\pi i/3}$ and $g$ acts as multiplication by $\omega$. In this case $H^{2}(W;T^{2})=H^{3}(W;\Lambda_{0})=\mathbb{Z}/3$. The split example occurs as a subgroup of the normalizer of the maximal torus inside $SU(3)$ (the permutation matrix for $(123)$ gives an order 3 lift of $g$ to $SU(3)$).
15.B. Lattices and Weyl groups
For subgroups we must classify $W$-invariant lattices in $\mathbb{Z}[\omega]$ and identify their Weyl groups. We have described the method in Subsection 4.B above. For this lattice the matrix $M$ is given by
$M=\left(\begin{array}[]{cc}-1&-1\\ 1&-2\end{array}\right)$
Lemma 15.1.
. (i) The unique rank 0 sublattice of $\Lambda^{0}$ is 0, and it corresponds to $S=T^{2}$. It has trivial Weyl group.
(ii) There are no rank 1 lattices.
Proof : There are no lines invariant under $g$ so every nonzero invariant lattice must be of rank 2. ∎
Lemma 15.2.
Each rank 2 lattice in $\Lambda^{0}=\mathbb{Z}[\omega]$ is generated as a $W$-module by a single element $a+b\omega$ which may be taken to have argument in $0,\pi/3)$. All these lattices are isomorphic to $\mathbb{Z}[\omega]$ as $W$-modules. The Weyl group is of order 3.
Proof : We choose a nonzero element $a+b\omega$ of $\Lambda$ of minimal length, and we may suppose the argument is in $0,\pi/3)$. This generates a lattice of equilateral triangles. If $\Lambda\neq\langle a+b\omega,\omega(a+b\omega)\rangle$ then any other element must be closer to a lattice point than $a+b\omega$ is to the origin. This contradicts the choice of $a+b\omega$. The order of the Weyl group follows from the fact that $\det(M)=3$.
15.C. Sections and conjugacy
Now we need understand the sequence
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
where $i^{S}_{*}$ is the focus of attention (we refer to Section 6 for a summary, including the cohomology groups).
Since there are no rank 1 sublattices, we turn directly to the case when $S$ is finite, and $\Lambda$ has rank 2. The lattices $\Lambda$ are all rank 1 free modules over $\mathbb{Z}[\omega]$, so they are isomorphic, with lattice duals isomorphic to $\mathbb{Z}[\omega]$. Thus $H^{i}(W;T/S)\cong H^{i}(W;T^{2})=H^{i+1}(W;\mathbb{Z}[\omega])$, which is zero if $i$ is odd and $\mathbb{Z}/3$ if $i>0$ is even. Thus lifts of $\epsilon(G)$ are unique up to conjugacy if they exist, and they exist if and only if $\epsilon(G)$ is annihilated by $a^{2}$. In the split case, there is always a lift, and if $\epsilon(G)\neq 0$ we may use the duality statements in Subsection 2.C to see that $\epsilon(G)$ lifts precisely when $\Lambda^{S}$ lies in $3\Lambda^{0}$.
15.D. Summary
The diagram is rather simple.
In the split case, it has one finite subgroup for each element of the set
$\mathcal{F}:=\{a+b\omega\in\mathbb{Z}[\omega]\;|\;\arg(a+b\omega)\in[0,\pi/3)\}$
together with a point for $G$. As a topological space it is the one point compactification of $\mathcal{F}$. The component structure has a group of order 3 over points of $\mathcal{F}$ and the trivial group over $G$.
In the non-split case, the diagram is exacly similar but with $\mathcal{F}$ replaced by the subspace
$\mathcal{F}_{0}:=\{a+b\omega\;|\;a+b\omega\in\mathcal{F},a\equiv b\equiv 0(3)\}.$
16. The case $W=D_{6}$, $\Lambda_{0}=\mathbb{Z}[\omega]^{\prime}$ or $\mathbb{Z}[\omega]^{\prime\prime}$
16.A. Preamble
We will deal with two different toral lattices in this section. Let $W=D_{6}$ and suppose the toral lattice is $\Lambda_{0}=\mathbb{Z}[\omega]^{\prime}$ or $\mathbb{Z}[\omega]^{\prime\prime}$ where the $2\pi/3$ rotation $\rho$ acts as multiplication by $\omega$ and the reflection $\tau$ in the $x$ axis acts by complex conjugation (ie reflection in the line through the number $v=1$) in the primed case and reflection in the line through $v=2+\omega$ in the double primed case. In this case $H^{i}(W;T^{2})=H^{i+1}(W;\Lambda_{0})=H^{i+1}(C_{3};\mathbb{Z}[\omega])^{C_{2}}$ . The relevant groups are
$H^{1}(\Sigma_{3};\mathbb{Z}[\omega]^{\prime})=\mathbb{Z}/3,H^{2}(\Sigma_{3};\mathbb{Z}[\omega]^{\prime})=0,H^{3}(\Sigma_{3};\mathbb{Z}[\omega]^{\prime})=0$
and
$H^{1}(\Sigma_{3};\mathbb{Z}[\omega]^{\prime\prime})=0,H^{2}(\Sigma_{3};\mathbb{Z}[\omega]^{\prime\prime})=0,H^{3}(\Sigma_{3};\mathbb{Z}[\omega]^{\prime\prime})=\mathbb{Z}/3.$
16.B. Realisation in $SU(3)$
The $\mathbb{Z}[\omega]^{\prime}$ case is realised as the normalizer of the maximal torus in $SU(3)$; an easy way to see this is by thinking first of $U(3)$ where $D_{6}\cong\Sigma_{3}$ acts by the permuting rows and columns.
We observe that the extension is split. The element $g$ realising $(123)$ is the ordinary permutation matrix as before, but the elements realizing the transpositions should be the negatives of the permutation matrices (since 3 is odd, this puts them in $SU(3)$). As discussed previously, the full subgroups are of the form $H(S,\sigma)$ where $\sigma$ is specified by an element $g$ of order 3 outside $T^{2}$.
16.C. Lattices and Weyl groups
For subgroups we must classify $W$-invariant lattices in $\mathbb{Z}[\omega]$ and identify their Weyl groups. We have described the method in Subsection 4.B above. In fact we have already classified the lattices invariant under $C_{3}$, so we need only restrict attention to those also invariant under the reflection in $\langle v\rangle$.
Lemma 16.1.
(i) The unique rank 0 sublattice of $\Lambda^{0}$ is 0, and it corresponds to $S=T^{2}$. It has trivial Weyl group.
(ii) There are no rank 1 lattices.
Proof : This follows from Lemma 17.1. ∎
Lemma 16.2.
Each rank 2 lattice in $\Lambda^{0}$ is generated as a $\mathbb{Z}[\omega]$-module by a single element $a+b\omega$ which is a multiple of $v$. The Weyl group is of order 3.
Proof : The minimal vectors are equally spaced at angles of $\pi/3$. Since they are preserved by reflection in the line of $v$, the line of $v$ must pass through one of the minimal vectors.
As for the Weyl group, the calculation for $C_{3}$ shows the order is at most 3, and we observe that the additional condition imposed by $\tau$ is automatically satisfied for these lattices, because one basis vector is $v$ and the other is $(1+\omega)v=(1+g)v$. ∎
16.D. Sections and conjugacy
Now we need understand the sequence
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
where $i^{S}_{*}$ is the focus of attention (we refer to Section 6 for a summary, including the cohomology groups).
We turn immediately to the case when $S$ is finite, so $\Lambda^{S}$ has rank 2. There are two cases.
Case 1: $v=1$.
In this case all lattices concerned are $\mathbb{Z}[\omega]^{\prime}$, so $H^{i}(W;\mathbb{T})=H^{i}(W;\mathbb{T}/S)=H^{i+1}(W;\mathbb{Z}[\omega]^{\prime})=0$ for $i=1,2$. In this case all extensions are split, and there is a single $G$-conjugacy class for each $S$.
Case 1: $v=2+\omega$.
The lattices $\Lambda$ are all copies of $\mathbb{Z}[\omega]^{\prime\prime}$ and $H^{i}(W;\mathbb{T})=H^{i}(W;\mathbb{T}/S)=H^{i+1}(W;\mathbb{Z}[\omega]^{\prime\prime})$ This is zero if $i=1$ and $\mathbb{Z}/3$ if $i=2$. Thus lifts of $\epsilon(G)$ are unique up to $G$-conjugacy if they exist, and they exist if and only if $\epsilon(G)$ is annihilated by $a^{2}$. As in Section 15, in the split case, there is always a lift, and if $\epsilon(G)\neq 0$ we may use the duality statements in Subsection 2.C to see that $\epsilon(G)$ lifts precisely when $\Lambda^{S}$ lies in $3\Lambda^{0}$.
16.E. Summary
There are three essentially different cases. In all three, the space $\mathrm{Sub}(G)/G$ is the one-point compactification of the space of finite subgroups.
If the module is $\mathbb{Z}[\omega]^{\prime}$ the extension is split and the finite full subgroups correspond to $\mathcal{F}=\{n\in\mathbb{Z}\;|\;n\geq 1\}$ where $n$ gives the lattice $\Lambda^{S}$ generated by $n$, which corresponds to $H=\mathbb{T}[n]\rtimes\Sigma_{3}$.
If the module is $\mathbb{Z}[\omega]^{\prime\prime}$ and the extension is split, the finite full subgroups correspond to $\mathcal{F}=\{n(2+\omega)\in\mathbb{Z}\;|\;n\geq 1\}$ where $n(2+\omega)$ gives the lattice $\Lambda^{S}$ generated by $n(2+\omega)$, which corresponds to $\mathbb{T}[n]\rtimes\Sigma_{3}$. In the non-split case $\mathcal{F}$ is replace by $\mathcal{F}_{0}$ of elements of $\mathcal{F}$ congruent to 0 mod 3.
17. The case $W=D_{12}$ acting on $\Lambda_{0}$ as a $G_{2}$-lattice
17.A. Preamble
This example is the normalizer of the maximal torus inside a group of local type $G_{2}$ and dimension 14. In other words $\Lambda_{0}$ is the $G_{2}$-lattice and taken with the Lie theoretic Weyl group $W=D_{12}$ acting on the lattice. In concrete terms $D_{12}$ acts as the symmetry group of $\Lambda_{0}=\mathbb{Z}[\omega]$. A generating reflection is complex conjugation, and a generating rotation is by $2\pi/6$. We will use $1$ and $\omega$ as our basis elements.
The group $D_{12}$ has $D_{6}$ as a normal subgroup. Since the cohomology of $D_{6}$ is 3-primary, we see
$H^{i}(D_{12};\mathbb{Z}[\omega])=H^{i}(D_{6};\mathbb{Z}[\omega]^{\prime})^{C_{2}}.$
This means $H^{i}(D_{12};\mathbb{Z}[\omega])=0$ for $i=2,3$, and the only group of this type is the split extension.
17.B. Lattices and Weyl groups
For subgroups we must classify $W$-invariant lattices in $\Lambda^{0}$ and identify their Weyl groups. We have described the method in Subsection 4.B above. For this lattice the group $W$ has two generators, and the two matrices $M$, for rotation and reflection respectively, are given by
$M_{1}=\left(\begin{array}[]{cc}0&-1\\ 1&-1\end{array}\right),M_{2}=\left(\begin{array}[]{cc}0&-1\\ 0&-2\end{array}\right)$
Lemma 17.1.
. (i) The unique rank 0 sublattice of $\Lambda^{0}$ is 0, and it corresponds to $S=T^{2}$. It has trivial Weyl group.
(ii) There are no rank 1 sublattices.
Lemma 17.2.
Every rank 2 lattice is a multiple of $\Lambda^{0}$ and has trivial Weyl group.
Proof : This follows since $\Sigma_{3}=D_{6}$ is a subgroup of $D_{12}$ and the lattice restricts to $\mathbb{Z}[\omega]^{\prime}$. ∎
17.C. Sections and conjugacy
Now we need understand the sequence
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
where $i^{S}_{*}$ is the focus of attention (we refer to Section 6 for a summary, including the cohomology groups).
The $W$-invariant lattices are all multiples of $\Lambda_{0}$, and all relevant cohomology groups are 0. There is thus one conjugacy class of full subgroups for each positive integer $m$.
17.D. Summary
The space of conjugacy classes of full subgroups is the one-point compactification of $\{1,2,3,\ldots\}$. The Weyl groups are all trivial and the subgroup corresponding to $m$ is $\mathbb{T}[m]\rtimes D_{12}$.